### v1

# Hypersurfaces with Large Automorphism Groups

Abstract.
We find sharp upper bounds on the order of the automorphism group of a hypersurface in complex projective space in every dimension and degree. In each case, we prove that the hypersurface realizing the upper bound is unique up to isomorphism and provide explicit generators for the automorphism group. Similar results were recently proven independently by S. Yang, X. Yu, and Z. Zhu using different methods [35].
2020 Mathematics Subject Classification:
14J50, 14J70
In this paper, we consider the following question for any positive integers $n,d$:
Question 0.1.
What is the largest possible automorphism group of a smooth hypersurface $X$ of degree $d$ in complex projective space $\mathbb{P}_{{\mathbb{C}}}^{n+1}$?
When $d$ is $1$ or $2$, all the smooth hypersurfaces for an arbitary fixed $n$ are isomorphic to one another and furthermore have infinite automorphism group. 0.1 is therefore meaningful only when $d\geq 3$. When $(n,d)=(1,3)$, $X$ is a genus $1$ curve and has an infinite automorphism group. When $(n,d)=(2,4)$, $X$ is a quartic K3 surface, and some of these also have infinite automorphism group (see, e.g. [24, Proof of Theorem 4]).
In all the remaining cases, $\mathrm{Aut}(X)$ is finite by results of Matsumura-Monsky [24].
For almost all values of $n,d$, the precise upper bound for $|\mathrm{Aut}(X)|$ was not known. Our main result is to calculate the upper bound exactly for every $n$ and $d$, completely answering 0.1.
Theorem 0.2.
Let $X_{d}\subset\mathbb{P}^{n+1}_{{\mathbb{C}}}$ be a smooth hypersurface of degree $d$, where $d\geq 3$ and $(n,d)\neq(1,3),(2,4)$. Then $|\mathrm{Aut}(X)|\leq(n+2)!d^{n+1}$ except for the following six cases, where the indicated bound holds:
  1. (1)
    $n=1,d=4$: $|\mathrm{Aut}(X)|\leq 168$,
  2. (2)
    $n=1,d=6$: $|\mathrm{Aut}(X)|\leq 360$,
  3. (3)
    $n=2,d=6$: $|\mathrm{Aut}(X)|\leq 6912$,
  4. (4)
    $n=2,d=12$: $|\mathrm{Aut(X)}|\leq 86400$,
  5. (5)
    $n=4,d=6$: $|\mathrm{Aut}(X)|\leq 6531840$,
  6. (6)
    $n=4,d=12$: $|\mathrm{Aut}(X)|\leq 186624000$.
Each of these bounds is sharp, and the hypersurface $X$ achieving the bound is unique up to isomorphism for every pair $(n,d)$.
The generic bound is achieved by the Fermat hypersurface. A stronger and more precise form of Theorem 0.2 is given in Theorem 2.3, where we explicitly describe the hypersurfaces and groups realizing each upper bound.
Right before this paper appeared, another work was announced by S. Yang, X. Yu, and Z. Zhu [35] with a different proof of Theorem 0.2. Both results use Collins’ works on the classification of subgroups of $\mathrm{GL}_{N}({\mathbb{C}})$ [6, 7], but the methods otherwise differ.
Outline. The outline of this paper is as follows. In Section 1, we introduce the necessary preliminaries in the representation theory of finite groups. Section 2 summarizes what was previously known and gives a more precise statement of Theorem 0.2, including equations and generators for the hypersurfaces with largest automorphism group in the exceptional cases. Section 3 introduces the machinery to prove the upper bounds in Theorem 0.2 and completes the proof in most dimensions. Section 4 finishes the proof in the remaining cases.
Acknowledgements. The authors would like to thank János Kollár for helpful discussions and comments.
1. Preliminaries
This section summarizes the material from group and representation theory that will be required for the proof of Theorem 0.2. Throughout, we will work over the complex numbers ${\mathbb{C}}$, so all representations of finite groups will be complex. All representations $V$ in the paper will be finite-dimensional.
1.1. Group Theory and Notations
We’ll use the following standard notations for various groups:
  • $\mu_{r}$, the cyclic group of order $r$,
  • $r\coloneqq\mu_{r}$, where no confusion arises,
  • $S_{r}$, the symmetric group on $r$ elements,
  • $A_{r}$, the alternating group on $r$ elements,
  • $H_{p^{m}}$, the Heisenberg group of unipotent $3\times 3$ matrices over ${\mathbb{F}}_{p}$ when $m=1$, or more generally an extraspecial group of order $p^{1+2m}$,
  • $\mathrm{GL}_{n}({\mathbb{F}})$, the general linear group of invertible $n\times n$ matrices over the field ${\mathbb{F}}$,
  • $\mathrm{SL}_{n}({\mathbb{F}})$, the special linear group of $n\times n$ matrices over ${\mathbb{F}}$ of unit determinant,
  • $\mathrm{PGL}_{n}({\mathbb{F}})$, the general projective linear group $\mathrm{GL}_{n}({\mathbb{F}})/F^{\times}$,
  • $\mathrm{Sp}_{n}({\mathbb{F}})$, the symplectic group of $n\times n$ matrices over ${\mathbb{F}}$ (with $n$ even), preserving a non-degenerate skew-symmetric bilinear form,
  • $\mathrm{PSp}_{n}({\mathbb{F}})$, the projective symplectic group of $n\times n$ matrices over ${\mathbb{F}}$.
The groups $\mathrm{SL}_{n}({\mathbb{F}})$, etc., are written $\mathrm{SL}_{n}(q)$, etc., when ${\mathbb{F}}={\mathbb{F}}_{q}$ is the finite field with $q$ elements.
There are a few other groups that appear in the paper. In particular, the simple groups $\mathrm{PSU}_{4}(3)$ and $O_{8}^{+}(2)$ of Lie type appear as Jordan-Hölder constituents of important subgroups of $\mathrm{GL}_{6}({\mathbb{C}})$ and $\mathrm{GL}_{8}({\mathbb{C}})$, respectively. Since there are simple descriptions for the matrix generators for these subgroups, we don’t provide additional definitions related to finite groups of Lie type, and instead refer to the references, e.g. [6].
Definition 1.1.
We say that $G$ is a extension of $\bar{G}$ by $H$ if there exists an exact sequence
$1\rightarrow H\rightarrow G\rightarrow\bar{G}\rightarrow 1$.
If $H$ is contained in $Z(G)$, then the extension is central.
We often use the notation of the group ATLAS to write down finite groups; for instance, the notation $3.A_{6}$ refers to the perfect extension of $A_{6}$ by the cyclic group $\mu_{3}$, also known as the Valentiner group.
Suppose that $\rho:G\rightarrow\mathrm{GL}(V)$ is a representation of the finite group $G$. We say that the representation $\rho$ is irreducible if it does not decompose nontrivially as a direct sum of representations. We frequently make use of the following consequence of Schur’s lemma:
Lemma 1.2.
Suppose $g\in Z(G)$ is an element of the center of $G$ and $\rho:G\rightarrow\mathrm{GL}(V)$ is an irreducible representation. Then $\rho(g)$ is a scalar matrix.
In particular, the center of an irreducible $G\subset\mathrm{GL}(V)$ coincides with its intersection with the scalar matrices.
We say that the representation $\rho$ as above is imprimitive if there is a nontrivial direct sum decomposition $V=V_{1}\oplus\cdots\oplus V_{r}$ such that the action of $G$ permutes $V_{1},\ldots,V_{r}$. That is, for each $g\in G$ and each $i=1,\ldots,r$, $g(V_{i})=V_{j}$ for some $j$. In particular, we allow imprimitive representations to be reducible.
A group representation $\rho$ is called primitive if it is not imprimitive. We call a subgroup $G$ of $\mathrm{GL}(V)$ primitive when the corresponding representation is so.
Given a representation $\rho:G\rightarrow\mathrm{GL}(V)$, there is an induced action of $G$ on the polynomial algebra over $V$ by $(g\cdot f)(\textbf{x})\coloneqq f(\rho(g)^{-1}\textbf{x})$. A polynomial $f$ is an invariant of the representation $\rho$ if $g\cdot f=f$ for each $g\in G$. We say $f$ is a semi-invariant if $g\cdot f=\chi(g)f$ for some linear character $\chi$ of $G$.
Lemma 1.3.
Let $\rho:G\rightarrow\mathrm{GL}(V)$ be a representation. Any semi-invariant polynomial of $\rho$ is an invariant polynomial of $\rho|_{N}$, where $N\subset G$ is a normal subgroup with $G/N$ abelian. In particular, if the group $G$ has no nontrivial abelian quotients, i.e., is perfect, then every semi-invariant polynomial of a representation of $G$ is invariant.
Conversely, if $N\subset G$ is a normal subgroup with $G/N$ finite abelian and there is an invariant polynomial of degree $d$ for $\rho|_{N}$, there is a semi-invariant polynomial of degree $d$ for $\rho$.
Proof.
Any semi-invariant polynomial $f$ satisfies $g\cdot f=\chi(g)f$, where $\chi$ is some one-dimensional representation of $G$. The polynomial $f$ is therefore invariant for the restriction of $\rho$ to the kernel $N$ of $\chi$. Since $G/N\subset{\mathbb{C}}^{*}$, it is abelian.
For the last statement, suppose $f$ of degree $d$ is invariant under $\rho|_{N}$. Then the span of the orbit of $f$ under the $G$-action is a finite-dimensional vector space of polynomials of degree $d$ with induced action by $G/N$. Since $G/N$ is abelian, there is an eigenvector for this action, which is a semi-invariant of degree $d$ for $\rho$. ∎
1.2. Primitive Subgroups of the General Linear Group
A result of Collins bounds the size of primitive subgroups of $\mathrm{GL}_{N}({\mathbb{C}})$ for every $N$. The proof of this result utilizes, at least in a small way, the classification of finite simple groups.
Theorem 1.4.
[6, Theorem A] Let $G\subset\mathrm{GL}_{N}({\mathbb{C}})$ be a primitive subgroup for $N\geq 2$. Then the index $[G:Z(G)]$ of the center of $G$ is bounded for fixed $N$. Moreover, if $[G:Z(G)]$ achieves this bound, $G/Z(G)\cong S_{N+1}$, unless $N=2,3,4,5,6,7,8,9,$ or $12$. In these exceptional cases, the bounds on $[G:Z(G)]$ are listed in Table 1. When the bound is achieved, $G/Z(G)\cong H/Z(H)$, where $H$ is the group shown in the third column.
$N$ Largest $[G:Z(G)]$ $H$
$2$ $60$ $2.A_{5}$
$3$ $360$ $3.A_{6}$
$4$ $25920$ $\mathrm{Sp}_{4}(3)$
$5$ $25920$ $\mathrm{PSp}_{4}(3)$
$6$ $6531840$ $6_{1}.\mathrm{PSU}_{4}(3).2_{2}$
$7$ $1451520$ $\mathrm{Sp}_{6}(2)$
$8$ $348364800$ $22.O^{+}_{8}(2).$2
$9$ $4199040$ $3^{1+4}.\mathrm{Sp}_{4}(3)$
$12$ $448345497600$ $6.\mathrm{Suz}$
Table 1. The exceptional primitive subgroups of $\mathrm{GL}_{N}(C)$. The second column gives the upper bound on $[G:Z(G)]$, while the third column gives the normal structure of a group $H$ with minimal $|Z(H)|$.
Remark 1.5.
[…]
Definition 1.6.
Let $\Xi(N)$ denote the function whose output is the upper bound on $[G:Z(G)]$ over all primitive subgroups $G$ of $\mathrm{GL}_{N}({\mathbb{C}})$. Then $\Xi(1)=1$ and $\Xi(N)=(N+1)!$ for each $N\geq 2$ apart from the exceptional values of $N$ from Table 1, at which $\Xi$ takes the corresponding value from the second column.
We also need some more specific information about the classification of primitive subgroups of $\mathrm{GL}_{N}({\mathbb{C}})$ for certain small values of $N$. Such groups have been completely classified for $N\leq 11$ by the work of many authors (see [6, Section 1] for some history). These results are frequently phrased as classifications of unimodular subgroups (i.e., subgroups of $\mathrm{SL}_{n}({\mathbb{C}})$). This distinction is relatively unimportant, by the following lemma:
Lemma 1.7.
Let $G\subset\mathrm{GL}(V)$ be a finite primitive subgroup. There is a finite primitive group $H\subset\mathrm{SL}(V)$ such that $G/Z(G)\cong H/Z(H)$ (so in particular, $G$ and $H$ have the same image in $\mathrm{PGL}_{N}({\mathbb{C}})$) and any polynomial semi-invariant for $G$ is a polynomial semi-invariant for $H$.
Proof.
We construct $H$ as follows. Let $g_{1},\ldots,g_{s}$ be a minimal generating set for $G$. For each $i=1,\ldots,s$, choose a scalar $a_{i}\in{\mathbb{C}}^{*}$ such that $\det(a_{i}g_{i})=1$. Then define: $H\coloneqq\langle a_{1}g_{1},\ldots,a_{s}g_{s}\rangle\subset\mathrm{SL}(V)$. The group $H$ is primitive (and in particular irreducible), because if there were subspaces $V_{1},\ldots,V_{r}\subset V$ permuted by $H$, then they would be permuted by each $a_{i}g_{i}$ and hence every $g_{i}$, contradicting the assumption that $G$ is primitive.
The centers $Z(G)$ and $Z(H)$ coincide with the scalar subgroups of $G$ and $H$ by Lemma 1.2. Since the images of $g_{i}$ and $a_{i}g_{i}$ are the same in $\mathrm{PGL}(V)$, the images $G/Z(G)$ and $H/Z(H)$ of $G$ and $H$, respectively, in this group are isomorphic. Finally, if $f$ is a polynomial semi-invariant of $G$, it follows from the definition of $H$ that $h\cdot f$ is a scalar multiple of $f$ for any $h\in H$. ∎
In light of Lemma 1.7, we can replace any primitive subgroup of $\mathrm{GL}_{N}({\mathbb{C}})$ with a unimodular subgroup having very similar properties. Since this in general replaces invariants with semi-invariants, we phrase all the results below in terms of semi-invariant polynomials.
Also, we note that some of the references (e.g., [11]) actually classify quasiprimitive subgroups of $\mathrm{GL}_{N}({\mathbb{C}})$, namely those for which the restriction of the representation to any normal subgroup is a direct sum of copies of exactly one irreducible representation of the subgroup. This is a weaker property than primitive (see, e.g., [32] for more), so the primitive groups are certainly included in the list. From now on, we won’t mention quasiprimitivity further; at worst, this could result in some unrealized cases in the lists.
The proof of Theorem 0.2 requires additional information about primitive subgroups of $\mathrm{GL}_{N}({\mathbb{C}})$ for $N=2,3,4,5,6,$ and $8$. All these cases besides $N=8$ are summarized in [12, Section 8.5]. For $N=2,3,4$, we also refer to the list of primitive subgroups with explicit matrix generators appears in [2], which can be used to compute invariant degrees in MAGMA or GAP.
Proposition 1.8.
Suppose $G$ is a primitive subgroup of $\mathrm{GL}_{2}({\mathbb{C}})$. Then $G/Z(G)\cong H/Z(H)$, where $H$ is one of the following:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
[…]
The smallest possible degree of a semi-invariant polynomial is $12$ in case (1), $6$ in case (2), and $4$ in case (3). In appropriate coordinates, the unique degree $12$ semi-invariant polynomial of $H=2.A_{5}$ may be written
$xy(x^{10}+11x^{5}y^{5}-x^{10}),$
and the unique degree $6$ semi-invariant polynomial of $H=2.S_{4}$ may be written
$x^{5}y-xy^{5}.$
Proof.
The classification of subgroups of $\mathrm{GL}_{2}({\mathbb{C}})$ is well-known and dates back to work of Klein in the 19th century [19]. The semi-invariant polynomials for the various groups are discussed in [30, Section 0.13] and [26, Section 1.2]. Their degrees correspond to geometric features of the corresponding polyhedron. In particular, the number of vertices of the icosahedron, octahedron, and tetrahedron are $12,6$, and $4$, respectively. ∎
Proposition 1.9.
Suppose $G$ is a primitive subgroup of $\mathrm{GL}_{3}({\mathbb{C}})$. Then $G/Z(G)\cong H/Z(H)$, where $H$ is one of the following:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
[…]
  4. (4)
[…]
The smallest possible degree of a semi-invariant polynomial is $6$ in case (1) and for the Hessian and its index $3$ subgroup in case (2). The smallest degree is $4$ in case (3). For case (4), the smallest degree semi-invariants are $2$ and $4$. In appropriate coordinates, the unique degree $4$ semi-invariant of the Klein group is the Klein quartic
$x^{3}y+y^{3}z+z^{3}x.$
The next smallest degree semi-invariant has degree $6$.
The unique degree $6$ semi-invariant of the Valentiner group is the Wiman sextic
$10x^{3}y^{3}+9(x^{5}+y^{5})z-45x^{2}y^{2}z^{2}-135xyz^{4}+27z^{6}.$
The next smallest degree semi-invariant has degree $12$.
Proof.
This classification dates back at least to [2, Chapter V], though Blichfeldt’s list “misses” the subgroups of $\mathrm{SL}_{3}({\mathbb{C}})$ which are the product of $\mathrm{PSL}_{2}(7)$ or $A_{5}$ with the scalar subgroup $\mu_{3}$. Blichfeldt’s work also includes explicit matrix generators.
The claims about semi-invariant polynomials of $3.A_{6}$ and $\mathrm{PSL}_{2}(7)$ are well-known and date back to Wiman [34] and Klein [18], respectively. The triple cover of the Hessian group, which has order $648$, is the extension of the nonabelian group $H_{3}$ of order $27$ and exponent $3$ by $\mathrm{SL}_{2}(3)\cong 2.A_{4}$. Its image in $\mathrm{PGL}_{3}({\mathbb{C}})$, the Hessian group, is the automorphism group of the Hesse configuration in $\mathbb{P}^{2}_{{\mathbb{C}}}$ consisting of $9$ points and $12$ lines. The smallest degree semi-invariant for this group is $6$ [1], and the same is true for the index $3$ primitive subgroup, by MAGMA computation. ∎
Proposition 1.10.
Suppose $G$ is a primitive subgroup of $\mathrm{GL}_{4}({\mathbb{C}})$ satisfying $[G:Z(G)]\geq 648$. Then $G/Z(G)\cong H/Z(H)$, where $H$ is one of the following:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    $H$ is an extension by a group of order at most $2$ of a group $A\times B/Z$, where $A,B$ are among (1), (2), (3) of Proposition 1.8 and $Z$ is a central subgroup of order $2$ contained in neither factor; here $[H:Z(H)]\leq 7200$;
  4. (4)
[…]
  5. (5)
[…]
The smallest degree for a polynomial semi-invariant is $12$ for (1) and $8$ for (4) and (5). The group $N$ in (2) fits into an exact sequence
[…]
where $\tilde{H}_{4}$ is generated by $H_{4}$ and $iI_{4}$. The smallest degree for a polynomial semi-invariant for subgroup in case (2) is always at least $4$, and is at least $8$ unless $H$ is the subgroup corresponding to the exceptional embedding of $S_{5}\subset S_{6}$ or its subgroup of index $2$. In this latter case, the unique semi-invariant polynomial of degree $4$ is
$x_{0}^{4}+x_{1}^{4}+x_{2}^{4}+x_{3}^{4}-6(x_{0}^{2}x_{1}^{2}+x_{0}^{2}x_{2}^{2}+x_{0}^{2}x_{3}^{2}+x_{1}^{2}x_{2}^{2}+x_{1}^{2}x_{3}^{2}+x_{2}^{2}x_{3}^{2}).$
[…]
Proof.
The classification of primitive subgroups of $\mathrm{GL}_{4}({\mathbb{C}})$ is originally due to Blichfeldt [2, Chapter VII]; all the groups above are listed (with generators) in his work. Note that the summary of the list in [12, Section 8.5] misses a few groups, including $2.A_{5}$ and some of the tensor product cases. Using the matrix generators, all the claims about semi-invariant polynomials may be verified in MAGMA. We note that the extension of $\tilde{H}_{4}$ by $S_{5}$ in case (2) with semi-invariant of degree $4$ is labeled “$19^{\circ}$” in [2, p. 172]. ∎
Proposition 1.11.
Suppose $G$ is a primitive subgroup of $\mathrm{GL}_{5}({\mathbb{C}})$ satisfying $[G:Z(G)]\geq 9720$. Then $G$ is generated by scalar matrices and the group $\mathrm{PSp}_{4}(3)\subset\mathrm{GL}_{5}({\mathbb{C}})$, which is simple of order $25920$. The smallest degree for a semi-invariant polynomial of this group is $4$.
Proof.
[…]
Proposition 1.12.
Suppose $G$ is a primitive subgroup of $\mathrm{GL}_{6}({\mathbb{C}})$ satisfying $[G:Z(G)]\geq 174960$. Then $G/Z(G)\cong H/Z(H)$, where $H$ is one of the following:
  1. (1)
[…]
  2. (2)
[…]
The smallest degree semi-invariant polynomial in case (2) is $12$, while in case (1) it is $6$; the unique semi-invariant polynomial of that degree is
$\sum_{i=0}^{5}x_{i}^{6}-10\sum_{0\leq i<j\leq 5}x_{i}^{3}x_{j}^{3}-180x_{0}x_{1}x_{2}x_{3}x_{4}x_{5}.$
Proof.
The classification in dimension $6$ is due to Lindsay [21], where we simply read off the size of each group. Note that his cases (I) and (II), which deal with tensor products of subgroups of $\mathrm{GL}_{2}({\mathbb{C}})$ and $\mathrm{GL}_{3}({\mathbb{C}})$ and subgroups thereof, have center of index at most $60\cdot 360=21600$, which is too small. The semi-invariants of $2.J_{2}$ are readily computable with the character tables in GAP, while a description of explicit matrix generators for $6_{1}.\mathrm{PSU}_{4}(3).2_{2}$ appears in [22, p. 407]. ∎
Proposition 1.13.
Suppose $G$ is a primitive subgroup of $\mathrm{GL}_{8}({\mathbb{C}})$ satisfying $[G:Z(G)]\geq 88179840$. Then $G/Z(G)\cong H/Z(H)$, where $H$ is one of the following:
  1. (1)
[…]
  2. (2)
    $H$ contains a subgroup isomorphic to a central product $T$ of an extraspecial group $H_{8}$ of order $2^{7}$ with $\mu_{4}$, and $H/T$ is a subgroup of $\mathrm{Sp}_{6}(2)$; such an $H$ has $[H:Z(H)]\leq 185794560$.
Moreover, any such $G$ has semi-invariant polynomials only in even degree.
Proof.
The classification in dimension $8$ is due to [16] when $7$ does not divide the order of the group and [11, Theorem A] in the remaining cases. Inspection of these lists reveals that only cases (iii) and (ix) of [11, Theorem A] have large enough order to be considered here; these are cases (2) and (1), respectively. Note that $22.O_{8}^{+}(2).$2 is also the Weyl group $W(E_{8})$, and a description of explicit matrix generators for this group appears in [22, p. 407]. In the case (iii),
$[H:Z(H)]\leq\frac{(2^{7}\cdot 4/2)|\mathrm{Sp}_{6}(2)|}{2}=185794560.$
Both of these cases yield only even degree semi-invariant polynomials because any abelian quotient of either contains $-I_{8}$. ∎
2. Background and Main Results
In this section, we’ll state a more precise version of Theorem 0.2, with a full description of the optimal examples in all the exceptional cases. We’ll also provide some background and motivation for this problem.
In order to study the automorphism group of the smooth hypersurface $X\coloneqq\{f=0\}\subset\mathbb{P}^{n+1}$ of degree $d$, we will focus on linear automorphisms of $X$, namely those which extend to the ambient variety $\mathbb{P}^{n+1}$. We denote by $\mathrm{Lin}(X)\subset\mathrm{Aut}(X)$ the subgroup of all linear automorphisms of $X$.
A well-known result is that $\mathrm{Aut}(X)=\mathrm{Lin}(X)$ whenever $n\geq 1$, $d\geq 3$, and $(n,d)$ is not $(1,3)$ or $(2,4)$. For $n\geq 3$, this is a consequence of the Grothendieck-Lefschetz theorem [13, Exposé XII, Corollaire 3.6], while the $n=2$ and $n=1$ cases are due to Matsumura-Monsky [24, Theorem 2] and Chang [5, Theorem 1]. The group $\mathrm{Lin}(X)$ is furthermore finite for $n\geq 1$ and $d\geq 3$ (this is stated for $n\geq 2$ in [24, Theorem 1], and the same proof works for plane curves). We’ll state our main result below in terms of $\mathrm{Lin}(X)$; this will imply Theorem 0.2 in light of the above comments, but is also stronger, since it includes the cases $(n,d)=(1,3),(2,4)$.
For most of the paper, we work with subgroups of $\mathrm{GL}_{N}({\mathbb{C}})$ instead of $\mathrm{PGL}_{N}({\mathbb{C}})$. Therefore, we define $\mathrm{Lin}(f)=\{g\in\mathrm{GL}_{n+2}({\mathbb{C}}):g\cdot f=f\}$.
Lemma 2.1.
The natural homomorphism $\mathrm{Lin}(f)\rightarrow\mathrm{Lin}(X)$ induced by $\mathrm{GL}_{n+2}({\mathbb{C}})\rightarrow\mathrm{PGL}_{n+2}({\mathbb{C}})$ is a central extension of $\mathrm{Lin}(X)$ by the group $\mu_{d}I$ of scalar matrices.
Proof.
For any representative $g\in\mathrm{GL}_{n+2}({\mathbb{C}})$ of an element in $\mathrm{Lin}(X)$, we must have $g\cdot f=af$ for some constant $a\in{\mathbb{C}}$. Composing $g$ with an appropriate scalar transformation yields an element in $\mathrm{Lin}(f)$ with the same image in $\mathrm{PGL}_{n+2}({\mathbb{C}})$, so $\mathrm{Lin}(f)\rightarrow\mathrm{Lin}(X)$ is surjective. Its kernel consists of the scalar matrices that preserve $f$, which is precisely $\mu_{d}I$ since $f$ is homogeneous of degree $d$. ∎
We’ll see that the example with largest linear automorphism group in most dimensions and degrees is the Fermat hypersurface.
Example 2.2.
For any positive integers $n$ and $d$, the Fermat hypersurface of dimension $n$ and degree $d$ is the hypersurface $X\subset\mathbb{P}^{n+1}$ defined by the equation
$x_{0}^{d}+x_{1}^{d}+\cdots+x_{n+1}^{d}=0.$
This hypersurface is smooth and it’s straightforward to see that its automorphism group contains the semidirect product $({\mathbb{Z}}/d)^{\oplus(n+1)}\rtimes S_{n+2}$, where $({\mathbb{Z}}/d)^{\oplus(n+1)}$ is the subgroup of automorphisms given by multiplying each coordinate by a root of unity (modulo overall scalar transformations) and $S_{n+2}$ permutes the coordinates. When $d\geq 3$ and $d-1$ is not a power of the characteristic of the field, we have $\mathrm{Lin}(X)=({\mathbb{Z}}/d)^{\oplus(n+1)}\rtimes S_{n+2}$ [28, 20]. In particular, $|\mathrm{Lin}(X)|=(n+2)!d^{n+1}$.
Similarly, $\mathrm{Lin}(f)=({\mathbb{Z}}/d)^{\oplus(n+2)}\rtimes S_{n+2}$, with order $|\mathrm{Lin}(f)|=(n+2)!d^{n+2}$.
Significant heuristic evidence suggests that the Fermat hypersurface should achieve the largest automorphism group, at least generically.
For instance, a result of Howard and Sommese [15] shows for each fixed $n$ that there exists a constant $C_{n}$ depending only on $n$ with the following property: For every smooth hypersurface $X$ of dimension $n$ and degree $d\geq 3$, $|\mathrm{Lin(X)}|\leq C_{n}d^{n+1}$. This shows at least that the Fermat hypersurface example has the right asymptotics in degree, though the result provides no effective procedure for computing $C_{n}$.
Also, a result of Collins [7] shows that the symmetric group $S_{N+1}$, embedded in $\mathrm{GL}_{N}({\mathbb{C}})$ via the regular representation, has the largest Jordan constant of any subgroup of $\mathrm{GL}_{N}({\mathbb{C}})$ for $N\geq 71$. The Jordan constant $J$ of the finite group $G$ is the smallest index $[G:N]$ of a normal abelian subgroup $N$ of $G$, and roughly measures “how far $G$ is from being abelian”. Since the symmetric group is asymptotically “largest” in this sense, it’s plausible that the Fermat hypersurface has largest automorphism group. Collins’ result doesn’t directly prove this, however, even for $N\geq 71$, and it furthermore suggests that exceptional examples of small dimension could exist. Our main result confirms these expectations.
Theorem 2.3.
Suppose that $X\subset\mathbb{P}^{n+1}_{{\mathbb{C}}}$ is a smooth hypersurface of degree $d$ with $n\geq 1$ and $d\geq 3$. The Fermat hypersurface, with linear automorphism group of order $(n+2)!d^{n+1}$, achieves the largest $|\mathrm{Lin}(X)|$ in dimension $n$ and degree $d$, except in the cases $(n,d)=(1,4),(1,6),(2,4),\\ (2,6),(2,12),(4,6),$ and $(4,12)$.
In each of these exceptional cases, the hypersurface with largest possible $|\mathrm{Lin}(X)|$ and the matrix generators for the linear automorphism group are described in Examples 2.5 through 2.11. For any $n,d$, the hypersurface with largest $|\mathrm{Lin}(X)|$ is unique up to isomorphism.
Remark 2.4.
  1. (1)
    A few cases of Theorem 2.3 were known previously. When $n=1$, the theorem was proved independently by works of Pambianco [29] and Harui [14], building off many previous results in specific degrees. Several recent works partially or completely classify the automorphism groups that can occur for other specific small values of $n$ and $d$. As a result, the Fermat hypersurface was already known to have the largest automorphism group for the following pairs $(n,d)$: $(2,3)$ [8], $(3,3)$ [33], $(3,5)$ [27], and $(4,3)$ [23]. All other cases were open, to the authors’ knowledge.
  2. (2)
    The case $(n,d)=(2,4)$ is not included in the original statement Theorem 0.2 because $\mathrm{Aut}(X)$ and $\mathrm{Lin}(X)$ can differ in this case and some quartic K3’s have infinite automorphism group.
We list the optimal hypersurfaces in the exceptional cases below, including matrix generators for $\mathrm{Lin}(f)$.
Example 2.5 ($n=1,d=4$).
Let $X\subset\mathbb{P}^{2}$ be the Klein quartic curve defined by
$f\coloneqq x_{0}^{3}x_{1}+x_{1}^{3}x_{2}+x_{2}^{3}x_{0}=0.$
The automorphism group of $X$ is $\mathrm{PSL}_{2}(7)$.
Explicitly, $\mathrm{Lin}(f)$ is an extension of $\mathrm{PSL}_{2}(7)$ by $\mu_{4}$ generated by the following matrices,
$M_{1}=i\begin{pmatrix}\epsilon^{4}&0&0\\ 0&\epsilon^{2}&0\\ 0&0&\epsilon\end{pmatrix},M_{2}=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&\epsilon\end{pmatrix},M_{3}=\alpha\begin{pmatrix}\epsilon-\epsilon^{6}&\epsilon^{2}-\epsilon^{5}&\epsilon^{4}-\epsilon^{3}\\ \epsilon^{2}-\epsilon^{5}&\epsilon^{4}-\epsilon^{3}&\epsilon-\epsilon^{6}\\ \epsilon^{4}-\epsilon^{3}&\epsilon-\epsilon^{6}&\epsilon^{2}-\epsilon^{5}\end{pmatrix},$
where $\epsilon$ is a primitive $7$th root of unity and $\alpha=\frac{1}{\sqrt{-7}}$.
This group has order $648$, and its quotient $\mathrm{Lin}(X)=\mathrm{PSL}_{2}(7)\subset\mathrm{PGL}_{3}({\mathbb{C}})$ has order $168$.
Example 2.6 ($n=1,d=6$).
Let $X\subset\mathbb{P}^{2}$ be the Wiman sextic curve defined by
$f\coloneqq 10x_{0}^{3}x_{1}^{3}+9(x_{0}^{5}+x_{1}^{5})x_{2}-45x_{0}^{2}x_{1}^{2}x_{2}^{2}-135x_{0}x_{1}x_{2}^{4}+27x_{2}^{6}=0.$
The automorphism group of $X$ is isomorphic to $A_{6}$. The above equation is the presentation of the Wiman sextic due to Wiman himself [34]. The matrix generators for the automorphism group in these coordinates are very messy. For completeness, we provide a set of matrix generators for $\mathrm{Lin}(f)$ in a different coordinate system:
$M_{1}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix},M_{2}=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix},M_{3}=\begin{pmatrix}1&0&0\\ 0&0&\zeta^{2}\\ 0&-\zeta&0\end{pmatrix},M_{4}=\frac{1}{2}\begin{pmatrix}1&\tau&\tau^{-1}\\ \tau^{-1}&\tau&1\\ \tau&-1&\tau^{-1}\end{pmatrix},$
where $\zeta$ is a primitive third root of unity and $\tau=\frac{1+\sqrt{5}}{2}.$
The group $\mathrm{Lin}(f)$ has order $2160$, and its quotient $\mathrm{Lin}(X)=A_{6}\subset\mathrm{PGL}_{3}({\mathbb{C}})$ has order $360$.
Example 2.7 ($n=2,d=4$).
Let $X\subset\mathbb{P}^{3}$ be the smooth surface defined by the equation
$f\coloneqq x_{0}^{4}+x_{1}^{4}+x_{2}^{4}+x_{3}^{4}-6(x_{0}^{2}x_{1}^{2}+x_{0}^{2}x_{2}^{2}+x_{0}^{2}x_{3}^{2}+x_{1}^{2}x_{2}^{2}+x_{1}^{2}x_{3}^{2}+x_{2}^{2}x_{3}^{2})=0.$
The linear automorphism group of $X$ fits into an exact sequence
$0\rightarrow\mu_{2}^{\oplus 4}\rightarrow\mathrm{Lin}(X)\rightarrow S_{5}\rightarrow 0,$
More explicitly, the group $\mathrm{Lin}(f)$ is generated by the matrices
$M_{1}=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{pmatrix},M_{2}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix},M_{3}=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&1\end{pmatrix},$
$M_{4}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{pmatrix},M_{5}=\frac{1+i}{2}\begin{pmatrix}-i&0&0&i\\ 0&1&1&0\\ 1&0&0&1\\ 0&-i&i&0\end{pmatrix},M_{6}=\frac{1+i}{\sqrt{2}}\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}.$
The subgroup of $\mathrm{GL}_{4}({\mathbb{C}})$ generated by these matrices has order $7680$ with scalar subgroup of order $4$. The image of this group in $\mathrm{PGL}_{4}({\mathbb{C}})=\mathrm{Aut}(\mathbb{P}^{3})$ has order $|\mathrm{Lin}(X)|=1920$.
Example 2.8 ($n=2,d=6$).
Let $X\subset\mathbb{P}^{3}_{{\mathbb{C}}}$ be the smooth surface defined by the equation
$f\coloneqq x_{0}^{5}x_{1}-x_{0}x_{1}^{5}+x_{2}^{5}x_{3}-x_{2}x_{3}^{5}=0.$
The automorphism group of $X$ is the wreath product of a central extension of $S_{4}$ by $\mu_{6}$ with $S_{2}$, modulo its center.
Explicitly, let
$M_{1}=\frac{1}{2}\begin{pmatrix}1+i&1+i\\ -1+i&1-i\end{pmatrix},M_{2}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},M_{3}=\begin{pmatrix}\epsilon&0\\ 0&\epsilon^{19}\end{pmatrix},$
where $\epsilon$ is a primitive $24$th root of unity with $\epsilon^{6}=i$.
The matrices $M_{1},M_{2},M_{3}$ generate a central extension of $S_{4}$ by $\mu_{6}$ in $\mathrm{GL}_{2}({\mathbb{C}})$. The linear automorphism group of $X$ is generated by $4\times 4$ block matrices of the form
$\begin{pmatrix}M_{i}&0\\ 0&M_{j}\end{pmatrix},$
for $i,j\in\{1,2,3\}$ and the block matrix
$\begin{pmatrix}0&I_{2}\\ I_{2}&0\end{pmatrix}.$
The subgroup of $\mathrm{GL}_{4}({\mathbb{C}})$ generated by these matrices has order $41472$, with scalar subgroup of order $6$. The image of this group in $\mathrm{PGL}_{4}({\mathbb{C}})=\mathrm{Aut}(\mathbb{P}^{3})$ has order $|\mathrm{Lin}(X)|=6912$.
Example 2.9 ($n=2,d=12$).
Let $X\subset\mathbb{P}^{3}_{{\mathbb{C}}}$ be the smooth surface defined by the equation
$f\coloneqq x_{0}x_{1}(x_{0}^{10}+11x_{0}^{5}x_{1}^{5}+x_{1}^{10})+x_{2}x_{3}(x_{2}^{10}+11x_{2}^{5}x_{3}^{5}+x_{3}^{10})=0.$
The automorphism group of $X$ is the wreath product of a central extension of $A_{5}$ by $\mu_{12}$ with $S_{2}$, modulo its center.
Explicitly, let
$M_{1}=\begin{pmatrix}\zeta&0\\ 0&\zeta^{49}\end{pmatrix},M_{2}=\frac{1}{\sqrt{5}}\begin{pmatrix}-\epsilon+\epsilon^{4}&\epsilon^{2}-\epsilon^{3}\\ \epsilon^{2}-\epsilon^{3}&\epsilon-\epsilon^{4},\end{pmatrix},$
where $\zeta$ is a primitive $60$th root of unity and $\epsilon=\zeta^{12}$.
The matrices $M_{1},M_{2}$ generate a central extension of $A_{5}$ by $\mu_{12}$ in $\mathrm{GL}_{2}({\mathbb{C}})$. The automorphism group of $X$ is generated by $4\times 4$ block matrices of the form
$\begin{pmatrix}M_{i}&0\\ 0&M_{j}\end{pmatrix},$
for $i,j\in\{1,2\}$ and the block matrix
$\begin{pmatrix}0&I_{2}\\ I_{2}&0\end{pmatrix}.$
The subgroup of $\mathrm{GL}_{4}({\mathbb{C}})$ generated by these matrices has order $1036800$, with scalar subgroup of order $12$. The image of this group in $\mathrm{PGL}_{4}({\mathbb{C}})=\mathrm{Aut}(\mathbb{P}^{3})$ has order $|\mathrm{Lin}(X)|=86400$.
Example 2.10 ($n=4,d=6$).
Let $X\subset\mathbb{P}^{5}_{{\mathbb{C}}}$ be the smooth hypersurface defined by the equation
$f\coloneqq\sum_{i=0}^{5}x_{i}^{6}-10\sum_{0\leq i<j\leq 5}x_{i}^{3}x_{j}^{3}-180x_{0}x_{1}x_{2}x_{3}x_{4}x_{5}=0.$
The automorphism group of $X$ is an extension of $\mathrm{PSU}_{4}(3)$ by $\mu_{2}$. Explicitly, this group is generated by the following matrices: all $6\times 6$ permutation matrices, all $6\times 6$ matrices of order $3$ and determinant $1$, and the matrix $I_{6}-Q/3$, where $Q$ has all entries equal to $1$.
The subgroup of $\mathrm{GL}_{6}({\mathbb{C}})$ generated by these matrices has order $39191040$, with scalar subgroup of order $6$. The image of this group in $\mathrm{PGL}_{6}({\mathbb{C}})=\mathrm{Aut}(\mathbb{P}^{5})$ has order $|\mathrm{Lin}(X)|=6531840$.
Example 2.11 ($n=4,d=12$).
Let $X\subset\mathbb{P}^{5}_{{\mathbb{C}}}$ be the smooth hypersurface defined by the equation
$x_{0}x_{1}(x_{0}^{10}+11x_{0}^{5}x_{1}^{5}+x_{1}^{10})+x_{2}x_{3}(x_{2}^{10}+11x_{2}^{5}x_{3}^{5}+x_{3}^{10})+x_{4}x_{5}(x_{4}^{10}+11x_{4}^{5}x_{5}^{5}+x_{5}^{10})=0.$
The automorphism group of $X$ is the wreath product of a central extension of $A_{5}$ by $\mu_{12}$ with $S_{3}$, modulo its center.
Explicitly, let
$M_{1}=\begin{pmatrix}\zeta&0\\ 0&\zeta^{49}\end{pmatrix},M_{2}=\frac{1}{\sqrt{5}}\begin{pmatrix}-\epsilon+\epsilon^{4}&\epsilon^{2}-\epsilon^{3}\\ \epsilon^{2}-\epsilon^{3}&\epsilon-\epsilon^{4},\end{pmatrix},$
where $\zeta$ is a primitive $60$th root of unity and $\epsilon=\zeta^{12}$.
The matrices $M_{1},M_{2}$ generate a central extension of $A_{5}$ by $\mu_{12}$ in $\mathrm{GL}_{2}({\mathbb{C}})$. The linear automorphism group of $X$ is generated by $6\times 6$ block matrices of the form
$\begin{pmatrix}M_{i}&0&0\\ 0&M_{j}&0\\ 0&0&M_{k}\end{pmatrix},$
for $i,j,k\in\{1,2\}$ and the block matrices
$\begin{pmatrix}0&I_{2}&0\\ I_{2}&0&0\\ 0&0&I_{2}\end{pmatrix},\begin{pmatrix}0&I_{2}&0\\ 0&0&I_{2}\\ I_{2}&0&0\end{pmatrix}.$
The subgroup of $\mathrm{GL}_{6}({\mathbb{C}})$ generated by these matrices has order $2239488000$, with scalar subgroup of order $12$. The image of this group in $\mathrm{PGL}_{4}({\mathbb{C}})=\mathrm{Aut}(\mathbb{P}^{5})$ has order $|\mathrm{Lin}(X)|=186624000$.
3. Bounding the Order of Automorphism Groups
In this section, we prove some general bounds on the automorphism groups of hypersurfaces. We work with the finite group $\mathrm{Lin}(f)$, the group of linear automorphisms preserving the equation $f$ of a hypersurface $X$. By Lemma 2.1, this only differs in order from the (linear) automorphism group of $X$ by a factor of the degree $d$.
3.1. A General Upper Bound Formula
Suppose $X\subset\mathbb{P}^{n+1}$ is a smooth hypersurface of degree $d$ with defining polynomial $f$. The idea for bounding $|\mathrm{Lin}(f)|$ is to decompose it as a subnormal series and bound each of the subquotients individually. To do this, we first introduce some notation relating to the group $G\coloneqq\mathrm{Lin}(f)$ that will be used for the next few sections.
We set $N\coloneqq n+2$ to be the dimension of the vector space $V\cong{\mathbb{C}}^{N}$ such that $\mathbb{P}(V)\cong\mathbb{P}^{n+1}$ is the ambient projective space for $X$. Thus we get a representation $G\subset\mathrm{GL}(V)$ of $G=\mathrm{Lin}(f)$. We can better understand this representation by decomposing it as follows:
Lemma 3.1.
For any subgroup $G\subset\mathrm{GL}(V)$, there is a direct sum decomposition $V=V_{1}\oplus\cdots\oplus V_{r}$ with the following properties:
  1. (1)
    $G$ permutes the subspaces $V_{1},\ldots,V_{r}$;
  2. (2)
    For each $i=1,\ldots,r$, the subgroup $H_{i}\coloneqq\mathrm{Stab}_{G}(V_{i})$ of elements sending $V_{i}$ to itself acts primitively on $V_{i}$.
This lemma is fairly standard. For completeness, we recall the proof, which also appears as [7, Lemma 1].
Proof.
If $G$ is primitive, the trivial decomposition satisfies the conditions of the lemma. Suppose instead that $V$ is an irreducible but not primitive $G$-representation. Then there exists an $H_{1}$-subrepresentation $V_{1}\subset V$ and a proper subgroup $H_{1}\subset G$ such that $V=V_{1}^{G}$ is induced from $V_{1}$. Choose $H_{1}$ a minimal subgroup with this property. Then $H_{1}$ acts primitively on $V_{1}$ and its images under $G$ form the remainder of the decomposition in the lemma. When $V$ is not irreducible, we can repeat this procedure for every irreducible subrepresentation. ∎
We call a direct sum $V=V_{1}\oplus\cdots\oplus V_{r}$ a primitive decomposition for $G$ if it satisfies the conditions of Lemma 3.1. This decomposition may not be unique.
From now on, fix a primitive decomposition $V=V_{1}\oplus\cdots\oplus V_{r}$ for $G=\mathrm{Lin}(f)$. This decomposition also has an associated dimension partition $\pi=(N^{\mu_{N}},\ldots,1^{\mu_{1}})$, where each $\mu_{k}$ is the number of times a subspace of dimension $k$ appears in the decomposition and $r=\mu_{N}+\cdots+\mu_{1}$.
We also set $H_{i}\coloneqq\mathrm{Stab}_{G}(V_{i})$, and let
$H\coloneqq\bigcap_{i=1}^{r}H_{i}$
be the kernel of the permutation action of $G$ on the spaces $V_{i}$, so that there is an embedding of $G/H$ in the symmetric group $S_{r}$. Then we can write the following subnormal series for $G$:
$1\trianglelefteq Z(H)\trianglelefteq H\trianglelefteq G,$
where $Z(H)$ is the center of $H$. The strategy for bounding $|G|$ will be to individually bound the orders of the subquotients $G/H$, $H/Z(H)$, and $Z(H)$. We also define $\hat{H}_{i}$ to be the image of $H_{i}$ in $\mathrm{GL}(V_{i})$ under the projection. This is called the $i$th primitive constituent of $G$.
The key to bounding the first two subquotients is a replacement theorem due to Collins that replaces $G$ with a better behaved subgroup $\hat{G}$ with the same primitive decomposition.
Theorem 3.2.
[7, Theorem 2] For any subgroup $G\subset\mathrm{GL}(V)$ with primitive decomposition $V=V_{1}\oplus\cdots\oplus V_{r}$, there exists a group $\hat{G}\subset\mathrm{GL}(V)$ with the same decomposition and such that, if we define
$\hat{H}=\bigcap_{i=1}^{r}\mathrm{Stab}_{\hat{G}}(V_{i}),$
then the following properties hold:
  1. (1)
    $H\subset\hat{H}$;
  2. (2)
    $\hat{H}=\hat{H}_{1}\times\cdots\times\hat{H}_{r}$, where each $\hat{H}_{i}$ acts faithfully and primitively on $V_{i}$ and trivially on all other direct summands; further, the image $\hat{H}_{i}\subset\mathrm{GL}(V_{i})$ coincides with the image of the projection $H_{i}\rightarrow\mathrm{GL}(V_{i})$;
  3. (3)
    $\hat{G}$ permutes the subgroups $\{\hat{H}_{i}\}$ under conjugation with the same permutation action as on the $\{V_{i}\}$ under the bijection $\hat{H}_{i}\mapsto V_{i}$;
  4. (4)
    $\hat{G}=\hat{H}\rtimes\hat{K}$, where $\hat{K}\subset S_{r}$ is the subgroup given by the direct product of symmetric groups acting on the orbits of the $\{V_{i}\}$ under $G$;
  5. (5)
    $[\hat{G}:\hat{H}]\geq[G:H]$, $[\hat{H}:Z(\hat{H})]\geq[H:Z(H)]$, and $Z(\hat{H})\supset Z(H)$.
A few of the properties listed in Theorem 3.2 are not in the statement of [7, Theorem 2] but are clearly stated in its proof.
The last ingredient we need is a method for bounding the center $Z(H)$. We first show that each matrix in $Z(H)$ is a block scalar matrix with respect to $V=V_{1}\oplus\cdots\oplus V_{r}$, meaning that it is contained in the direct sum of scalar subgroups of the $\mathrm{GL}(V_{i})$. Indeed, any matrix in $Z(\hat{H})$ is block scalar since the representation of each factor $\hat{H}_{i}$ in $\hat{H}$ on $V_{i}$ is primitive, in particular irreducible, so the conclusion follows from Lemma 1.2. Since $Z(\hat{H})\supset Z(H)$ by Theorem 3.2, every matrix in $Z(H)$ is also block scalar with respect to the primitive decomposition. Conversely, any block scalar matrix commutes with every element of $H$ because $H\subset\mathrm{GL}(V_{1})\oplus\cdots\oplus\mathrm{GL}(V_{r})$, so $Z(H)$ equals the block scalar subgroup of $H$.
Now we can bound $|Z(H)|$ using a generalization of a result of the first author [10, Theorem 3.3, Step 3], which bounded the order of an abelian subgroup of $\mathrm{Lin}(f)$:
Theorem 3.3.
Suppose $X=\{f=0\}\subset\mathbb{P}^{n+1}=\mathbb{P}(V)$ is a smooth hypersurface of degree $d\geq 3$. Let $V=V_{1}\oplus\cdots\oplus V_{r}$ be any decomposition of $V$ into a direct sum of positive-dimensional subspaces. Then the intersection of $\mathrm{Lin}(f)$ with the group of block scalar matrices with respect to this decomposition has order at most $d^{r}$.
Note that the decomposition here is arbitrary, and doesn’t actually need to be a primitive decomposition for $\mathrm{Lin}(f)$. Since any abelian subgroup is diagonalizable (hence block scalar for some decomposition into one-dimensional subspaces), [10, Theorem 3.3, Step 3] is a special case of Theorem 3.3 for $r=N=n+2$. The fact from Lemma 2.1 that the intersection of $\mathrm{Lin}(f)$ with the scalar matrices is $\mu_{d}$ is the special case $r=1$.
Proof.
The only consequence of the smoothness of $X$ that we need for the proof is the existence of certain monomials with nonzero coefficients in $f$. In particular, for each variable $x_{j}$, where $j=0,\ldots,n+1$, $f$ contains a monomial of the form $x_{j}^{d}$ or $x_{j}^{d-1}x_{k}$, $k\neq j$, with nonzero coefficient in $f$. Otherwise, $X$ would be singular at the coordinate point of the variable $x_{j}$ ([10, Proposition 1.2]).
Suppose first that each subspace of the decomposition is $1$-dimensional, so that we’re bounding the size of the diagonal subgroup $A$ of $\mathrm{Lin}(f)$. This case was proven in [10, Theorem 3.3, Step 3], but we review the proof for the reader’s convenience. Each diagonal automorphism $x_{j}\mapsto c_{j}x_{j}$ in $A$ must preserve each monomial of $f$ individually, so if
$f=\sum_{i=1}^{s}K_{i}\prod_{j=0}^{n+1}x_{j}^{m_{ij}}$
is invariant, then for each $i$,
(1) $\prod_{j=0}^{n+1}c_{j}^{m_{ij}}=1.$
Assemble the exponents $m_{ij}$ into a $s\times(n+2)$ matrix $M$. The equations above for each $i$ imply that $(\log|c_{0}|,\ldots,\log|c_{n+1}|)^{\mathsf{T}}\in\ker(M)$. We claim that the kernel of $M$ is trivial. To see this, take the $(n+2)\times(n+2)$ square minor $B$ of $M$ which corresponds to the monomials of the special form $x_{j}^{d}$ or $x_{j}^{d-1}x_{k}$, where the monomial for $x_{j}$ goes in the $j$th row. Then $B$ is invertible and in particular has positive determinant bounded above by $d^{n+2}$ [10, Lemma 3.6]. It follows that $\ker(M)=0$ and each $c_{j}$ is a unit complex number $c_{j}=e^{2\pi i\theta_{j}}$. The condition (1) then gives that $M(\theta_{0},\ldots,\theta_{n+1})^{\mathsf{T}}\in{\mathbb{Z}}^{s}$. The weaker condition $B(\theta_{0},\ldots,\theta_{n+1})^{\mathsf{T}}\in{\mathbb{Z}}^{n+2}$ is already enough to show that $(\theta_{0},\ldots,\theta_{n+1})$ is in the superlattice of ${\mathbb{Z}}^{n+2}$ spanned by $B^{-1}e_{0},\ldots,B^{-1}e_{n+1}$, where $e_{0},\ldots,e_{n+1}$ are the standard basis vectors of ${\mathbb{Z}}^{n+2}$. Since each $\theta_{j}$ is only defined modulo ${\mathbb{Z}}$, we have an injection $A\subset B^{-1}{\mathbb{Z}}^{n+2}/{\mathbb{Z}}^{n+2}$. This quotient has order $\det(B)\leq d^{n+2}$, completing the argument.
Now suppose $V=V_{1}\oplus\cdots\oplus V_{r}$ is an arbitrary decomposition and let $A$ be the intersection of $\mathrm{Lin}(f)$ with the block scalar matrices of this decomposition. We can define a linear transformation $V\rightarrow{\mathbb{C}}^{r}$ by “collapsing” the coordinates in each block. Indeed, if $x_{0},\ldots,x_{\dim(V_{1})-1}$ are the coordinates of the first block, map $x_{0}\mapsto\beta_{0}y_{1},\ldots,x_{\dim(V_{1})-1}\mapsto\beta_{\dim(V_{1})-1}y_{1}$, where $\beta_{0},\ldots,\beta_{\dim(V_{1})-1}$ are nonzero constants. After doing similarly on the other blocks, $V\rightarrow{\mathbb{C}}^{r}$ is $A$-equivariant for a diagonal action of $A$ on ${\mathbb{C}}^{r}$. Furthermore, the transformed polynomial $\bar{f}(y_{1},\ldots,y_{m})$ is $A$-invariant. It’s not necessarily true that $\bar{f}$ defines a smooth hypersurface any longer, but for a general choice of constants $\beta$, $\bar{f}$ will still contain monomials of the form $y_{j}^{d}$ or $y_{j}^{d-1}y_{k}$ with nonzero coefficients for each $j=1,\ldots,r$. This is because monomials of this type collapse to monomials of this type, and different monomials will not cancel for general $\beta$. Now, using the same proof as above, we have $|A|\leq d^{r}$, as required. ∎
We are now ready to prove a general bound on the order of $\mathrm{Lin}(f)$:
Theorem 3.4.
Using the notation from the beginning of Section 3.1:
$|\mathrm{Lin}(f)|\leq\mu_{1}!\mu_{2}!\cdots\mu_{N}!\prod_{i=1}^{r}[\hat{H}_{i}:Z(\hat{H}_{i})]d^{r}\leq\mu_{1}!\mu_{2}!\cdots\mu_{N}!\prod_{i=1}^{r}\Xi(\dim(V_{i}))d^{r}.$
More precisely,
$|Z(H)|\leq d^{r},$
$[H:Z(H)]\leq\prod_{i=1}^{r}[\hat{H}_{i}:Z(\hat{H}_{i})]\leq\prod_{i=1}^{r}\Xi(\dim(V_{i})),\text{and }$
$[G:H]\leq\mu_{1}!\mu_{2}!\cdots\mu_{N}!.$
Recall that the $\Xi$ function has values which are the upper bounds for $[G:Z(G)]$ over primitive linear groups $G$ in a particular dimension (see Definition 1.6).
Proof.
[…]
Next, $[H:Z(H)]\leq[\hat{H}:Z(\hat{H})]$ by Theorem 3.2, and $[\hat{H}:Z(\hat{H})]=\prod_{i=1}^{r}[\hat{H}_{i}:Z(\hat{H}_{i})]$, which is the expression that appears on the right of the first inequality. Since each $\hat{H}_{i}$ acts primitively and faithfully on $V_{i}$, the quantity $[\hat{H}_{i}:Z(\hat{H}_{i})]$ is bounded above by $\Xi(\dim(V_{i}))$ by definition.
Finally, we know from Theorem 3.2 that $[G:H]\leq[\hat{G}:\hat{H}]$, and $\hat{K}=\hat{G}/\hat{H}$ is the subgroup of $S_{r}$ given by the direct product of symmetric groups of orbits of the $\{V_{i}\}$ under $G$. Of course, the order of this group is bounded above by $r!$, but we can do better: it is in fact a subgroup of $S_{\mu_{N}}\times\cdots\times S_{\mu_{1}}\subset S_{r}$, where each factor acts only on the subset of $\{V_{i}\}$ of subspaces of the given dimension. This is because a subspace of dimension $k$ can only be permuted by $G$ to a subspace of the same dimension. Hence $[G:H]\leq\mu_{1}!\cdots\mu_{N}!$, completing the proof. ∎
3.2. Exceptional Partitions
Notice that the last expression in Theorem 3.4 depends only on the partition $\pi$ of $N$ and the degree $d$. We define
$B(\pi,d)\coloneqq\mu_{1}!\mu_{2}!\cdots\mu_{N}!\prod_{i=1}^{r}\Xi(\dim(V_{i}))d^{r}.$
An important special case of the above is $\pi=(1^{N})$. When $X$ is a Fermat hypersurface, $G$ has this $\pi$ as the associated partition and $G=\mathrm{Lin}(f)$ realizes the equality $|G|=N!d^{N}=B(\pi,d)$.
For two partitions $\pi$ and $\pi^{\prime}$ of the integers $N$ and $N^{\prime}$, respectively, the concatenation $\pi\oplus\pi^{\prime}$ of $\pi$ and $\pi^{\prime}$ is the partition of $N+N^{\prime}$ whose set of blocks is given by the disjoint union of sets of blocks of $\pi$ and $\pi^{\prime}$. The following lemma summarizes a few simple properties of the function $B(\pi,d)$:
Lemma 3.5.
Let $\pi$ be a partition of $N$ and $\pi^{\prime}$ a partition of $N^{\prime}$. Then:
  1. (1)
    $B(\pi,d)B(\pi^{\prime},d)\leq B(\pi\oplus\pi^{\prime},d)$, and
  2. (2)
    If the sets of block sizes of $\pi$ and $\pi^{\prime}$ are disjoint, $B(\pi,d)B(\pi^{\prime},d)=B(\pi\oplus\pi^{\prime},d)$.
Proof.
The term $d^{r}\prod_{i=1}^{r}\Xi(\dim(V_{i}))$ of $B(\pi,d)$ is clearly multiplicative for partition concatenation, so the inequality (1) follows from noticing that $a!b!\leq(a+b)!$ for any nonnegative integers $a$ and $b$. In the case that $\pi$ has no block of the same size as any block of $\pi^{\prime}$, the number of blocks of any given size is 00 for either $\pi$ or $\pi^{\prime}$ and the inequality of factorials becomes an equality. ∎
Using this bound, we can already prove that the Fermat hypersurface always has the largest automorphism group of any hypersurface if the dimension is large enough:
Theorem 3.6.
Fix a positive integer $N\geq 27$ and a degree $d\geq 3$. Then $B(\pi,d)<B((1^{N}),d)$ for any partition $\pi$ of $N$ such that $\pi\neq(1^{N})$.
When $X=\{f=0\}\subset\mathbb{P}^{n+1}$ is the Fermat hypersurface, $\mathrm{Lin}(f)=B((1^{N}),d)$, so Theorem 3.6 already shows that the Fermat hypersurface achieves the maximum order automorphism group in every degree $d\geq 3$ and dimension $n=N-2\geq 25$.
Proof.
We continue to use the notation from Theorem 3.4. That proposition shows that
$|\mathrm{Lin}(f)|\leq B(\pi,d).$
Notice that the right-hand side depends only on the degree and the partition $\pi$ of $N$. The Fermat hypersurface corresponds to the partition $\pi=(1^{N})$ and realizes the inequality $|\mathrm{Lin}(f)|=N!d^{N}$. We need to characterize the cases when
$B(\pi,d)=\mu_{1}!\mu_{2}!\cdots\mu_{N}!\prod_{i=1}^{r}\Xi(\dim(V_{i}))d^{r}>N!d^{N}=B((1^{N}),d).$
The strategy is to show that replacing blocks of $\pi$ with blocks of dimension $1$ always increases the quantity $B(\pi,d)$ apart from a small collection of cases. Indeed, for a given partition $\pi$ of $N$, we can write $\pi=\pi_{1}\oplus\cdots\oplus\pi_{N}$, where the partition $\pi_{i}$ contains only the blocks of size $i$. By Lemma 3.5(2), $B(\pi,d)=B(\pi_{1},d)\cdots B(\pi_{N},d)$. We will successively replace the partitions $\pi_{i}$ with partitions with blocks of smaller sizes in such a way that the modified $B(\pi,d)$ is larger. Indeed, if $\tau$ is a refinement of $\pi_{i}$, $\pi^{\prime}$ is the new partition of $N$ with $\pi_{i}$ replaced by $\tau$, and $B(\pi_{i},d)\leq B(\tau,d)$, then
$B(\pi,d)=B(\pi_{1},d)\cdots B(\pi_{N},d)\leq B(\pi,d)\cdots B(\tau,d)\cdots B(\pi_{N},d)\leq B(\pi^{\prime},d).$
This argument requires justifying many inequalities involving the function $B(\pi,d)$. Note that in each modification, we replace all the blocks of a given size with blocks of strictly smaller size. In each of these inequalities, there are more blocks in the new partition than the old, and thus $B(\pi,d)$ has a higher power of $d$ on the right. This allows us to simplify matters and assume that $d$ takes the smallest value $d=3$. Also, nearly all the inequalities contain a parameter $k$ or $\ell$. In each one, it’s straightforward to justify the inequality by verifying the base case where the parameter takes the smallest indicated value and then using induction.
We begin by noting that $k!(i+1)!3^{k}<(ki)!3^{ki}=B((1^{ki}),3)$ holds for all $k\geq 1$, $i>2$. The left-hand side of this inequality equals $B((i^{k}),3)$ unless $i$ is one of the special values from Theorem 1.4. Therefore, if $\pi$ contains $k$ blocks of size $i\neq 2,3,4,5,6,7,8,9$, or $12$, we modify $\pi$ by replacing all blocks of size $i$ by blocks of size $1$. By the above, this modification increases $B(\pi,d)$. For $i=12,9,7$, we claim that performing the same modification increases $B(\pi,d)$. This is justified by the following inequalities, which hold for all positive integers $k$:
$B((12^{k}),3)=k!448345497600^{k}3^{k}<(12k)!3^{12k}=B((1^{12k}),3);$
$B((9^{k}),3)=k!4199040^{k}3^{k}<(9k)!3^{9k}=B((1^{9k}),3);$
$B((7^{k}),3)=k!1451520^{k}3^{k}<(7k)!3^{7k}=B((1^{7k}),3).$
Next, for $i=8,6,5$, we claim that it’s always advantageous to replace each block of size $i$ with blocks of size $2$, plus one block of size $1$ if $i$ is odd. This is justified by the following inequalities, which hold for all positive integers $k$:
$B((8^{k}),3)=k!348364800^{k}3^{k}<(4k)!60^{4k}3^{4k}=B((2^{4k}),3);$
$B((6^{k}),3)=k!6531840^{k}3^{k}<(3k)!60^{3k}3^{3k}=B((2^{3k}),3);$
$B((5^{k}),3)=k!25920^{k}3^{k}<k!(2k)!60^{2k}3^{3k}=B((2^{2k},1^{k}),3).$
This leaves us with a partition $\pi$ containing only blocks of sizes $1,2,3$, and $4$. We can perform various additional reductions as follows. By the inequalities below for the indicated values of $k$ or $\ell$, we can replace the blocks of sizes $4$ and $3$ with blocks of sizes $2$ and $1$ unless there is at most one block of size $4$ or $3$.
$B((4^{k}),3)=k!25920^{k}3^{k}<(2k)!60^{2k}3^{2k}=B((2^{2k}),3),k\geq 2;$
$B((3^{2\ell}),3)=(2\ell)!360^{2\ell}3^{2\ell}<(3\ell)!60^{3\ell}3^{3\ell}=B((2^{3\ell}),3),\ell\geq 1;$
$B((3^{2\ell+1}),3)=(2\ell+1)!360^{2\ell+1}3^{2\ell+1}<(3\ell+1)!60^{3\ell+1}3^{3\ell+2}=B((2^{3\ell+1},1),3),\ell\geq 1.$
At this point $\pi$ only has blocks of sizes $4,3,2,1$, with at most one each of sizes $4$ and $3$. If $N\geq 8$, we can eliminate any lone remaining $4$ and $3$ blocks, due to the following inequalities, and the fact that $\pi$ would have to contain a subpartition of one of the following types: $(4,3),(4,2^{k}),k\geq 1,(4,1^{k}),k\geq 4,(3,2^{k}),k\geq 2,(3,1^{k}),k\geq 2$. Moreover, we can arrange so that this subpartition contains all blocks of $\pi$ of the indicated sizes, so that Lemma 3.5(2) applies.
$B((4,3),3)=25920\cdot 360\cdot 3^{2}<3!60^{3}3^{4}=B((2^{3},1),3);$
$B((4,2^{k}),3)=k!25920\cdot 60^{k}3^{k+1}<(k+2)!60^{k+2}3^{k+2}=B((2^{k+2}),3),k\geq 1;$
$B((4,1^{k}),3)=k!25920\cdot 3^{k+1}<(k+4)!3^{k+4}=B((1^{k+4}),3),k\geq 4;$
$B((3,2^{k}),3)=k!360\cdot 60^{k}3^{k+1}<(k+1)!60^{k+1}3^{k+2}=B((2^{k+1},1),3),k\geq 2;$
$B((3,1^{k}),3)=k!360\cdot 3^{k+1}<(k+3)!3^{k+3}=B((1^{k+3}),3),k\geq 2.$
Finally, under the assumption that $N\geq 8$, we’ve now reached the point that any $\pi$ has been replaced by a new $\pi$ with a larger $B(\pi,d)$ such that $\pi$ contains only $1$’s and $2$’s. The last step is to show that if $N$ is large enough, we can replace all the $2$’s with $1$’s. Note that if the number of $2$’s is at least $14$, we can replace all $2$’s with $1$’s:
$B((2^{k}),3)=k!60^{k}3^{k}<(2k)!3^{2k}=B((1^{2k}),3),k\geq 14.$
This means we can arrange for $\pi$ to have at most $13$ $2$’s. For each $1\leq k\leq 13$, one can verify:
$B((2^{k},1^{\ell}),3)=\ell!k!60^{k}3^{k+\ell}<(2k+\ell)!3^{2k+\ell}=B((1^{2k+\ell}),3),\ell\geq 27-2k.$
For this last inequality, one takes $\ell=27-2k$ as the base case and argues by induction on $\ell$ for each fixed $k$. In summary, we’ve shown that for $N\geq 27$, $B(\pi,d)$ attains its maximum value among partitions $\pi$ of $N$ when $\pi=(1^{N})$ for any $d\geq 3$. ∎
The hypersurface $X=\{f=0\}$ of degree $d\geq 3$ can only have an automorphism group of order equal to or larger than that of the Fermat hypersurface of the same dimension and degree if $G\coloneqq\mathrm{Lin}(f)$ has an associated partition $\pi$ with $B(\pi,d)\geq B((1^{N}),d)$, where $B((1^{N}),d)$ is the order of the automorphism group of the degree $d$ Fermat hypersurface. We say that such a $\pi\neq(1^{N})$ is exceptional for degree $d$. In light of Theorem 3.6, $\pi$ must be a partition of $N\leq 26$. A computer search of all such partitions yields exactly $80$ with the property that $B(\pi,3)\geq B((1^{N}),3)$ and $\pi\neq(1^{N})$; we simply call these exceptional partitions. They represent the remaining cases that need to be checked for large automorphism groups. Table 2 lists these partitions, along with the maximum $d$ for which each is exceptional, and the ratio $B/B_{F}\coloneqq B(\pi,d)/B((1^{N}),d)$.
Number $N$ Partition Max $d$ $B/B_{F}$
No. 1 2 $(2)$ 30 10.0
No. 2 3 $(3)$ 7 6.66
No. 3 3 $(2,1)$ 10 3.33
No. 4 4 $(4)$ 10 40.0
No. 5 4 $(3,1)$ 3 1.67
No. 6 4 $(2^{2})$ 17 33.3
No. 7 4 $(2,1^{2})$ 5 1.67
No. 8 5 $(5)$ 3 2.67
No. 9 5 $(4,1)$ 6 8.00
No. 10 5 $(3,2)$ 5 6.67
No. 11 5 $(2^{2},1)$ 7 6.67
No. 12 5 $(2,1^{3})$ 3 1.00
No. 13 6 $(6)$ 6 37.3
No. 14 6 $(4,2)$ 6 26.7
No. 15 6 $(4,1^{2})$ 4 2.67
No. 16 6 $(3^{2})$ 4 4.45
No. 17 6 $(3,2,1)$ 3 1.11
No. 18 6 $(2^{3})$ 12 66.7
No. 19 6 $(2^{2},1^{2})$ 4 2.22
No. 20 7 $(6,1)$ 4 5.33
No. 21 7 $(5,2)$ 3 1.27
No. 22 7 $(4,3)$ 4 7.62
No. 23 7 $(4,2,1)$ 4 3.81
No. 24 7 $(4,1^{3})$ 3 1.14
No. 25 7 $(3,2^{2})$ 4 6.35
No. 26 7 $(2^{3},1)$ 6 9.53
No. 27 8 $(8)$ 3 3.95
No. 28 8 $(6,2)$ 4 13.3
No. 29 8 $(6,1^{2})$ 3 1.33
No. 30 8 $(4^{2})$ 5 45.7
No. 31 8 $(4,2^{2})$ 5 19.0
No. 32 8 $(3^{2},2)$ 3 1.59
No. 33 8 $(2^{4})$ 9 95.2
No. 34 8 $(2^{3},1^{2})$ 4 2.38
No. 35 9 $(6,3)$ 3 2.96
No. 36 9 $(6,2,1)$ 3 1.48
No. 37 9 $(4^{2},1)$ 3 5.08
No. 38 9 $(4,3,2)$ 3 2.12
No. 39 9 $(4,2^{2},1)$ 3 2.12
No. 40 9 $(3^{3})$ 3 1.06
Number $N$ Partition Max $d$ $B/B_{\operatorname{F}}$
No. 41 9 $(3,2^{3})$ 4 5.29
No. 42 9 $(2^{4},1)$ 5 10.6
No. 43 10 $(6,4)$ 3 7.11
No. 44 10 $(6,2^{2})$ 3 5.93
No. 45 10 $(4^{2},2)$ 4 10.2
No. 46 10 $(4^{2},1^{2})$ 3 1.02
No. 47 10 $(4,2^{3})$ 4 12.7
No. 48 10 $(2^{5})$ 7 106
No. 49 10 $(2^{4},1^{2})$ 3 2.12
No. 50 11 $(4^{2},3)$ 3 1.85
No. 51 11 $(4,2^{3},1)$ 3 1.15
No. 52 11 $(3,2^{4})$ 3 3.85
No. 53 11 $(2^{5},1)$ 4 9.62
No. 54 12 $(6^{2})$ 3 3.02
No. 55 12 $(6,4,2)$ 3 1.08
No. 56 12 $(6,2^{3})$ 3 2.70
No. 57 12 $(4^{3})$ 3 11.1
No. 58 12 $(4^{2},2^{2})$ 3 3.08
No. 59 12 $(4,2^{4})$ 4 7.70
No. 60 12 $(2^{6})$ 6 96.2
No. 61 12 $(2^{5},1^{2})$ 3 1.60
No. 62 13 $(3,2^{5})$ 3 2.47
No. 63 13 $(2^{6},1)$ 4 7.40
No. 64 14 $(6,2^{4})$ 3 1.18
No. 65 14 $(4^{3},2)$ 3 1.22
No. 66 14 $(4^{2},2^{3})$ 3 1.01
No. 67 14 $(4,2^{5})$ 3 4.23
No. 68 14 $(2^{7})$ 5 74.0
No. 69 14 $(2^{6},1^{2})$ 3 1.06
No. 70 15 $(3,2^{6})$ 3 1.41
No. 71 15 $(2^{7},1)$ 3 4.93
No. 72 16 $(4,2^{6})$ 3 2.11
No. 73 16 $(2^{8})$ 4 49.3
No. 74 17 $(2^{8},1)$ 3 2.90
No. 75 18 $(2^{9})$ 4 29.0
No. 76 19 $(2^{9},1)$ 3 1.53
No. 77 20 $(2^{10})$ 3 15.3
No. 78 22 $(2^{11})$ 3 7.27
No. 79 24 $(2^{12})$ 3 3.16
No. 80 26 $(2^{13})$ 3 1.27
Table 2. The $80$ exceptional partitions of a positive integer $N$ for which $B(\pi,d)\geq B((1^{N}),d)$ for some $d\geq 3$. The fourth column shows the largest integer $d$ for which this occurs, and the fifth shows the value $B(\pi,3)/B((1^{N}),3)$ to three significant figures.
4. Finishing the Proof of Main Theorem
From the results of the last section, it only remains to analyze the finite number of cases listed in Table 2, all of which correspond to dimensions $n=N-2$ at most $24$.
Though there are $80$ exceptional partitions, a vast majority of these will not be realized by hypersurfaces with automorphism group larger than or equal to that of the Fermat. This is because the analysis in Section 3.2 does not use any information about invariant theory, which further constrains the degrees $d$ that can occur for various partitions. In this section, we’ll finish the proof by considering each of the exceptional partitions.
One subtlety is that the exceptional partitions each have some “wiggle room”: one need not take the largest possible group associated to a given $\pi$ to beat the Fermat automorphism group; instead, any group with partition $\pi$ whose order is within the factor $B/B_{F}\coloneqq B(\pi,3)/B((1^{N}),3)$ of the upper bound $B(\pi,3)$ is a priori a candidate for $\mathrm{Lin}(f)$ for some $d=\deg(f)\geq 3$. Therefore, we have to be careful when eliminating various partitions.
We’ll continue to use the notation from the beginning of Section 3.1. So we choose a primitive decomposition $V=V_{1}\oplus\cdots\oplus V_{r}$ for $G\coloneqq\mathrm{Lin}(f)$ as in Lemma 3.1, set $H_{i}\coloneqq\mathrm{Stab}_{G}(V_{i})$, and define $\hat{H}_{i}$ to be the image of $H_{i}$ in $\mathrm{GL}(V_{i})$. We call $\hat{H}_{i}\subset\mathrm{GL}(V_{i})$ the $i$th primitive constituent of $G$. As above, we also let
$H\coloneqq\bigcap_{i=1}^{r}H_{i}.$
Note that we are using the original group $G$ here rather than its replacement $\hat{G}$ from Theorem 3.2; replacing the group can change its invariant polynomials. Nevertheless, by Theorem 3.2(2), the image of $H_{i}$ in $\mathrm{GL}(V_{i})$ coincides with the $\hat{H}_{i}$ that appears in the replacement group. Finally, we’ll introduce the notation $\bar{H}_{i}$ for the image of $\hat{H}_{i}$ in $\mathrm{PGL}(V_{i})$. Since $Z(H)$ coincides with the block scalar subgroup of $H$, we have an induced inclusion
(2) $H/Z(H)\xhookrightarrow{\phi}\bar{H}_{1}\times\cdots\times\bar{H}_{r}\subset\mathrm{PGL}(V_{1})\oplus\cdots\oplus\mathrm{PGL}(V_{r}).$
induced by $H\hookrightarrow\hat{H}_{1}\times\cdots\times\hat{H}_{r}$.
The idea of this section is as follows: the upper bound $B(\pi,d)$ for $|G|$ is achieved only when each of the three subquotients of $G$ from Theorem 3.4 attains the upper bound for its order. Conversely, if the order of any subquotient falls short by some amount, $|G|$ will fall short of $B(\pi,d)$ by at least the same amount. In particular, if the bound is achieved, the inclusion in (2) must be an isomorphism. We use the following argument throughout this section, so we list it here for reference:
Setup 4.1.
For some $i$ among $1,\ldots,r$, consider the composition
$H/Z(H)\xhookrightarrow{\phi}\bar{H}_{1}\times\cdots\times\bar{H}_{r}\rightarrow\bar{H}_{1}\times\cdots\times\widehat{\bar{H}_{i}}\times\cdots\times\bar{H}_{r}.$
(The second map is the projection to the product omitting $\bar{H}_{i}$.) The kernel $\bar{K}$ of this composition can be identified with a subgroup of $\bar{H}_{i}$. Denote by $K$ its preimage in $H$. As a subgroup of $\mathrm{GL}(V_{1})\oplus\cdots\oplus\mathrm{GL}(V_{r})$, $K$ maps to a block scalar subgroup under the projection forgetting $\mathrm{GL}(V_{i})$, because the image of each element of $K$ in the projective linear groups of the other factors is trivial. Hence the kernel $K^{\prime}$ of this projection map is a subgroup of $H\subset G$ such that (1) $K^{\prime}$ is a normal subgroup of $K$ with $K/K^{\prime}$ abelian; (2) $K^{\prime}$ acts trivially on $V_{j}$, where $j\neq i$; and (3) the image $\bar{K}^{\prime}$ of $K^{\prime}\rightarrow\mathrm{PGL}(V_{i})$ is a normal subgroup of $\bar{K}$ with abelian quotient.
Since $f$ is an invariant polynomial of $G$, it is an invariant polynomial of $K^{\prime}$. The group $K^{\prime}$ acts only on $V_{i}$ and trivially on other summands, so setting the variables outside $V_{i}$ to be arbitrary constants, the resulting polynomial $f$ in the variables of $V_{i}$ must always be an invariant of the representation $K^{\prime}\rightarrow\mathrm{GL}(V_{i})$. If $d_{0}$ is the minimum positive degree for such an invariant, $f$ must have degree $d\geq d_{0}$, because it has degree at least $d_{0}$ in the variables corresponding to $V_{i}$ (these variables must appear somewhere in the polynomial because $f$ defines a smooth hypersurface). Finally, by Lemma 1.3, we can conclude that the same $d_{0}$ above occurs as the degree of a semi-invariant for $\bar{K}$ because $\bar{K}/\bar{K}^{\prime}$ is abelian. (To be clear, by semi-invariant for $\bar{K}\subset\mathrm{PGL}(V_{i})$, we mean a semi-invariant of any scalar extension of $\bar{K}$ in $\mathrm{GL}(V_{i})$.) In summary, $d\geq d_{0}$, where $d_{0}$ is the minimum positive degree of a semi-invariant of $\bar{K}$.
We will refer to the results from 4.1 throughout the section and use the notation from it, whenever there is no possibility for confusion.
The following lemma will also be useful for finding monomials excluding certain variables.
Lemma 4.2.
Let $X=\{f=0\}\subset\mathbb{P}^{n+1}$ be a smooth hypersurface of degree $d>1$. Then there is a monomial in $f$ with nonzero coefficient which does not involve any given set of $k$ variables, for any $k<\frac{n+2}{2}$.
Note that $\frac{n+2}{2}=\frac{N}{2}$, where $N$ is the dimension of the associated vector space $V$.
Proof.
Suppose every monomial in $f$ includes at least one of a given set of $k$ variables, say $x_{0},\ldots,x_{k-1}$. Then $X$ contains the $(n-k+1)$-plane $\{x_{0}=\cdots=x_{k-1}=0\}\subset\mathbb{P}^{n+1}$. But then $n-k+1\leq\frac{n}{2}$ by [9, Corollary 6.26] because $X$ is smooth. Rearranging gives $k\geq\frac{n+2}{2}$. Thus, when $k<\frac{n+2}{2}$, there is a monomial in $f$ not containing any of the $k$ variables. ∎
4.1. Preliminary Reductions
This subsection is a series of lemmas that eliminate all partitions for which exceptional hypersurface examples do not occur. First, we prove a result about the degree $d$ when $\pi$ contains at least one block of size $2$.
Lemma 4.3.
Suppose $G$ has a partition $\pi$ with $\mu_{2}\geq 1$. Then
  1. (1)
    If $d<12$:
    $\frac{B(\pi,d)}{|G|}\geq\min\left\{60,\left(\frac{5}{2}\right)^{\mu_{2}}\right\}.$
  2. (2)
    If $d=3$ and $4\mu_{2}>N$, we get further that
    $\frac{B(\pi,d)}{|G|}\geq\min\left\{120,12\left(\frac{5}{2}\right)^{\mu_{2}},2\cdot 5^{\mu_{2}}\right\}.$
Proof.
First, we’ll prove item (1). We consider two cases. The first is when a central extension of $A_{5}$ occurs as at least one of the primitive constituents. Without loss of generality, set $\bar{H}_{1}=A_{5}$. Consider 4.1 for $i=1$. The kernel $\bar{K}$ of the composition is either trivial or all of $A_{5}$ since $A_{5}$ is simple. If the kernel is trivial, then $H/Z(H)$ has index at least $60$ in $\bar{H}_{1}\times\cdots\times\bar{H}_{r}$, so $\frac{B(\pi,d)}{|G|}\geq 60$, so the condition in (1) is satisfied.
If $\bar{K}=A_{5}$, then we conclude by 4.1 and Proposition 1.8 that $d\geq 12$, contradicting the assumption $d<12$ in (1).
The second case is that a central extension of $A_{5}$ does not appear among the primitive constituents. This means that for each $V_{i}$ of dimension $2$, $|\bar{H}_{i}|\leq|S_{4}|=24$ by the classification in Proposition 1.8. This reduces the maximum possible order of $H/Z(H)$ by a factor of at least
$\left(\frac{60}{24}\right)^{\mu_{2}}=\left(\frac{5}{2}\right)^{\mu_{2}},$
completing the proof.
Now we prove (2). Since $4\mu_{2}>N$, it follows from Lemma 4.2 that there are monomials in $f$ with nonzero coefficient and variables only from the spaces $V_{i}$ of dimension $2$. Apply Theorem 3.2 to replace $G$ with a new group $\hat{G}$. Writing $\hat{H}$ as in Theorem 3.2, we know $Z(H)$ is a subgroup of $Z(\hat{H})=Z(\hat{H}_{1})\times\cdots\times Z(\hat{H}_{n})$. (In this product, each factor $Z(\hat{H}_{i})$ acts trivially on all other $V_{j},j\neq i$.) We claim that $Z(H)$ cannot equal $Z(\hat{H})$, just by virtue of the degree being odd. Indeed, all central extensions of $A_{4},S_{4}$, or $S_{5}$ have even order centers so the matrix $-I_{V}\subset Z(\hat{H})$. Hence if $Z(H)=Z(\hat{H})$, then since $f$ is invariant under $Z(H)$, the monomials involving only these variables from dimension $2$ summands $V_{i}$ must have even degree, hence $f$ has even degree, a contradiction. Thus, $|Z(H)|$ falls short of the maximum $|Z(\hat{H})|\leq d^{r}$ by a factor of at least $2$, because $Z(H)$ is a proper subgroup of $Z(\hat{H})$.
Now we again consider the possible primitive constituents. We already know from the proof of (1) in the lemma that $[H:Z(H)]$ falls short of the maximum by at least $60$ if $A_{5}$ is among the $\bar{H}_{i}$. Therefore, suppose it’s not. If some primitive constituent is a central extension of $S_{4}$, without loss of generality suppose $\bar{H}_{1}=S_{4}$. Again consider 4.1 for $i=1$. The kernel $\bar{K}$ of the composition is then either $S_{4}$, $A_{4}$, the Klein four-subgroup of $S_{4}$, or trivial. If it $S_{4}$ of $A_{4}$, $d\geq 4$ by 4.1 and Proposition 1.8, a contradiction. Therefore, $\bar{K}$ has order at most $4$ and $[H:Z(H)]$ is less than the product of the $|\bar{H}_{i}|$ by at least a factor of $24/4=6$, which in turn is less than the maximum theoretical value of $[H:Z(H)]$ by at least $\left(\frac{5}{2}\right)^{\mu_{2}}$, since no $A_{5}$ constituents appear. Finally, if there are no primitive constituents which are $A_{5}$ or $S_{4}$, we have that the possible order $[H:Z(H)]$ is reduced by
$\left(\frac{60}{12}\right)^{\mu_{2}}=5^{\mu_{2}}.$
In summary, $[H:Z(H)]$ falls short of its maximum from Theorem 3.4 by at least $60$, at least $6\cdot\left(\frac{5}{2}\right)^{\mu_{2}}$, or at least $5^{\mu_{2}}$, depending on the case. Since we showed $[Z(\hat{H}):Z(H)]\geq 2$ also, we can multiply these bounds to obtain the statement in the lemma. ∎
We can use this lemma to eliminate most of the candidate partitions in Table 2. For clarity, we’ll divide into two corollaries, each of which uses one of the parts of the lemma.
Corollary 4.4.
Suppose that $G=\mathrm{Lin}(f)$ has an exceptional partition which is one of the following: No. 7, No. 12, No. 17, No. 19, No. 21, No. 26, No. 32, No. 34, No. 36, No. 38-39, No. 41-42, No. 44, No. 47, No. 49, No. 51-53, No. 55-56, No. 58-59, No. 61-67, No. 69-80. Then $|G|<|B((1^{N}),d)|$.
Proof.
In each listed case, the maximum degree listed in Table 2 is less than $12$ and
$B(\pi,3)/B((1^{N}),3)$ is less than $\min\left\{60,\left(\frac{5}{2}\right)^{\mu_{2}}\right\}$ so the result holds by Lemma 4.3(1). ∎
Corollary 4.5.
Suppose that $G=\mathrm{Lin}(f)$ has an exceptional partition which is one of the following: No. 11, No. 25, No. 33, No. 48, No. 60, No. 68. Then $|G|<|B((1^{N}),d)|$.
Proof.
Each of these partitions satisfies $4\mu_{2}>N$ so Lemma 4.3(2) applies when $d=3$. Since $B(\pi,3)/B((1^{N}),3)$ is less than the minimum in the lemma, an exceptional example cannot occur for $d=3$. Thus we can assume $d\geq 4$ and we can replace the quotient $B(\pi,3)/B((1^{N}),3)$ in the table with $B(\pi,4)/B((1^{N}),4)$ to apply the reasoning of Lemma 4.3(1). We verify that this quotient is smaller than $\min\left\{60,\left(\frac{5}{2}\right)^{\mu_{2}}\right\}$ in each case. ∎
Next we can eliminate some partitions with blocks of size $3$.
Lemma 4.6.
Suppose $G$ has partition $\pi$ with $\mu_{3}\geq 1$. Then if $d<6$:
$\frac{B(\pi,d)}{|G|}\geq\min\left\{5,\left(\frac{15}{7}\right)^{\mu_{3}}\right\}.$
Proof.
Arrange the composition $V=V_{1}\oplus\cdots\oplus V_{r}$ so that $\dim(V_{1})=3$. We consider two cases. The first case is when the first primitive constituent $\hat{H}_{1}$ has quotient $\bar{H}_{1}$ equal to $A_{6}$ or the Hessian group $({\mathbb{Z}}/3)^{\oplus 2}\rtimes\mathrm{SL}_{2}(3)$ of order $216$. Consider 4.1 for $i=1$. The kernel $\bar{K}$ of the composition there is a normal subgroup of $\bar{H}_{1}$. Since $A_{6}$ is simple and the largest proper normal subgroup of $({\mathbb{Z}}/3)^{\oplus 2}\rtimes\mathrm{SL}_{2}(3)$ is $({\mathbb{Z}}/3)^{\oplus 2}\rtimes Q_{8}$, $\bar{K}$ either all of $\bar{H}_{1}$ or has order at most $216/3=72$. In the latter situation, we get that $|H/Z(H)|$ falls short of the maximum possible order by at least $360/72=5$, so the condition of the lemma is satisfied.
If the kernel is all of $\bar{H}_{1}$, we apply the conclusion of 4.1 and the facts about semi-invariant degrees from Proposition 1.9 to conclude $d\geq 6$, contradicting the assumption.
Finally, if neither of these types of primitive constituents appear among the $\hat{H}_{i}$, the largest possible constituent for a block of size $3$ is a central extension of $\mathrm{PSL}_{2}(7)$. In this case, the maximum possible order of $H/Z(H)$ is reduced by a factor of at least
$\left(\frac{360}{168}\right)^{\mu_{3}}=\left(\frac{15}{7}\right)^{\mu_{3}},$
completing the proof of the lemma. ∎
Corollary 4.7.
Suppose that $G=\mathrm{Lin}(f)$ has an exceptional partition which is one of the following: No. 5, No. 16, No. 40, No. 50. Then $|G|<|B((1^{N}),d)|$.
Proof.
In each listed case, the maximum degree listed in Table 2 is less than $6$ and
$B(\pi,3)/B((1^{N}),3)$ is less than $\min\left\{5,\left(\frac{15}{7}\right)^{\mu_{3}}\right\}$. ∎
Next, we can eliminate some partitions with blocks of size $4$.
Lemma 4.8.
Suppose $G$ has partition $\pi$ with $\mu_{4}\geq 1$. Then if $d<8$,
$\frac{B(\pi,d)}{|G|}\geq\min\left\{720,\left(\frac{18}{5}\right)^{\mu_{4}}\right\}$
Proof.
As usual, there are two cases. For the first, suppose there is a primitive constituent (which we can assume is $\hat{H}_{1}$) that falls into case (1) or (2) of Proposition 1.10, where in the second case we assume $\hat{H}_{1}/Z(\hat{H}_{1})=N/Z(N)$ for the full normalizer $N$ of $H_{4}\subset\mathrm{GL}_{4}({\mathbb{C}})$. Now consider 4.1 for $i=1$. If $\bar{H}_{1}=\mathrm{PSp}_{4}(3)$, which is simple, the kernel $\bar{K}$ is the whole group or trivial. If $\bar{H}_{1}=N/Z(N)$ instead, we note that any normal subgroup of $\bar{H}_{1}$ is index $2$ (corresponding to $A_{6}\subset S_{6}$) or has index at least $720$. So either $[H:Z(H)]$ falls short of its maximum by at least $720$, or $\bar{K}$ is among $\mathrm{PSp}_{4}(3),N/Z(N)$, or the index $2$ normal subgroup of $N/Z(N)$. In this latter situation, by 4.1 and Proposition 1.10, $d\geq 8$, a contradiction.
The second case is that each constituent is not of the type above, so that by Proposition 1.10, $[\hat{H}_{i}:Z(\hat{H}_{i})]\leq 7200$ for each $i$. This means that the order of $[H:Z(H)]$ falls short of the maximum possible by at least
$\left(\frac{25920}{7200}\right)^{\mu_{4}}=\left(\frac{18}{5}\right)^{\mu_{4}}.$
This finishes the lemma. ∎
Corollary 4.9.
Suppose that $G=\mathrm{Lin}(f)$ has an exceptional partition which is one of the following: No. 15, No. 24, No. 37, No. 45-46, No. 57. Then $|G|<|B((1^{N}),d)|$.
Proof.
In each listed case, the maximum degree listed in Table 2 is less than $8$ and
$B(\pi,3)/B((1^{N}),3)$ is less than $\min\left\{720,\left(\frac{18}{5}\right)^{\mu_{4}}\right\}$ so the result holds by Lemma 4.8. ∎
Lemma 4.10.
Suppose $G$ has partition $\pi$ with $\mu_{6}\geq 1$. Then if $d<6$,
$\frac{B(\pi,d)}{|G|}\geq\min\left\{604800,\left(\frac{112}{3}\right)^{\mu_{6}}\right\}.$
Proof.
Arrange the summands so $\dim(V_{1})=6$. There are two cases, depending on what $\hat{H}_{1}$ is. In the first case, $\hat{H}_{1}$ is one of the groups listed in Proposition 1.12. In this event, we consider 4.1 for $i=1$. The possibilities for $\bar{K}$ are $\mathrm{PSU}_{4}(3).2_{2}$, its index $2$ normal subgroup $\mathrm{PSU}_{4}(3)$, the Janko group $J_{2}$, or the trivial group. If $\bar{K}$ is trivial, $H/Z(H)$ has index at least $|J_{2}|=604800$ in the product of the $\hat{H}_{i}$, giving the first bound in the lemma. For other possible $\bar{K}$, we conclude from 4.1 and Proposition 1.12 that $d\geq 6$, a contradiction.
The second possibility is that the constituents above do not appear, so by Proposition 1.12, the maximum possible $[H:Z(H)]$ is at most $174960$. Therefore, the maximum possible order of $H/Z(H)$ is reduced from its theoretical upper bound in Theorem 3.4 by a factor of at least
$\left(\frac{6531840}{174960}\right)^{\mu_{6}}=\left(\frac{112}{3}\right)^{\mu_{6}}.$
Corollary 4.11.
Suppose that $G=\mathrm{Lin}(f)$ has an exceptional partition which is one of the following: No. 20, No. 28-29, No. 35, No. 43, No. 54. Then $|G|<|B((1^{N}),d)|$.
Proof.
In each listed case, the maximum degree listed in Table 2 is less than $6$ and
$B(\pi,3)/B((1^{N}),3)$ is less than $\min\left\{604800,\left(\frac{112}{3}\right)^{\mu_{6}}\right\}$ so the result holds by Lemma 4.10. ∎
At this point, we only have a handful of partitions left to eliminate. These will be treated individually.
Lemma 4.12.
Suppose that $G=\mathrm{Lin}(f)$ has an exceptional partition which is one of the following: No. 3, No. 8-10, No. 14, No. 22-23, No. 27, No. 30-31. Then $|G|<|B((1^{N}),d)|$.
Proof.
For each partition, we assume $|G|\geq B((1^{N}),d)$ and derive a contradiction.
Partitions No. 8: $\pi=(5)$ and No. 27: $\pi=(8)$.
In these cases, $G$ is primitive. In order for $|G|\geq B((1^{N}),d)$ for some $d\geq 3$ we need $[G:Z(G)]d\geq N!d^{N}$ so $[G:Z(G)]\geq N!d^{N-1}$. Even for $d=3$, the right hand side is $97205!\cdot 3^{4}=$ 9720 or $881798408!\cdot 3^{7}=$ 88179840, precisely the bounds that appear in Proposition 1.11 and Proposition 1.13, respectively. But the maximum possible degree from Table 2 is $3$ in both cases and a semi-invariant polynomial of degree at least $3$ must actually satisfy $d\geq 4$ for all the primitive groups listed in Proposition 1.11 and Proposition 1.13.
The next few cases make use of Lemma 4.2.
Partition No. 3: $\pi=(2,1)$. By Lemma 4.2, some monomials in $f$ involve only the variables from $V_{1}$ of dimension $2$, so $f(x_{0},x_{1},0)$ is nonzero and invariant under the primitive representation $\hat{H}_{1}\subset GL(V_{1})$. Since $d\leq 10$ by Table 2, $\bar{H}_{1}$ cannot be $A_{5}$. If it is $S_{4}$, $d\geq 6$ by Proposition 1.8 but the maximum order of $|G|$ from Theorem 3.4 is $24\cdot 6^{2}<6\cdot 6^{3}=B((1^{3}),6)$. If it is $A_{4}$, $d\geq 4$ but $12\cdot 4^{2}<6\cdot 4^{3}=B((1^{3}),4)$.
Partition No. 9: $\pi=(4,1)$. By Lemma 4.2, some monomials in $f$ involve only the variables from $V_{1}$ of dimension $4$. Hence $f(x_{0},x_{1},x_{2},x_{3},0)$ is nonzero and invariant under the primitive representation $\hat{H}_{1}\subset\mathrm{GL}(V_{1})$. Since $B(\pi,3)/B((1^{5}),3)=8$, $[\hat{H}_{1}:Z(\hat{H}_{1})]$ is within a factor of $8$ of the largest possible value, $25920$. By Proposition 1.10, no group with this property has a semi-invariant of degree at most $6$ (the max degree from the table), except for powers of the quadric $x_{0}x_{3}-x_{1}x_{2}$ in case (3). If $d=4$ or $6$ and $f(x_{0},x_{1},x_{2},x_{3},0)$ is a power of that quadric, it would mean a smooth threefold hypersurface of degree $4$ or $6$ would contain a multiple quadric as a hyperplane section, contradicting the Grothendieck-Lefschetz theorem, which says ${\operatorname{Pic}}(\mathbb{P}^{4})\rightarrow{\operatorname{Pic}}(X)$ is an isomorphism.
Partition No. 10: $\pi=(3,2)$. Let $\dim(V_{1})=3$ and $\dim(V_{2})=2$. Consider 4.1 for $i=2$. If $\bar{H}_{2}$ is $A_{5}$, the kernel $\bar{K}$ is either all of $A_{5}$ or trivial. If it is trivial, $[H:Z(H)]$ is reduced by a factor of $60$ from its maximum value, and we’re done since the value $B/B_{F}$ in Table 2 is 6.66. Otherwise, $d\geq 12$, which is greater than the max degree from the table, a contradiction.
Hence we’ve shown $|\bar{H}_{2}|\leq 24$, and Theorem 3.4 gives $|G|\leq 24[\hat{H}_{1}:Z(\hat{H}_{1})]d^{2}$. Comparing this to the Fermat hypersurface, we need $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 135$ and $d=3$ or $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 320$ and $d=4$. By Lemma 4.2, the polynomial $f(x_{0},x_{1},x_{2},0,0)$ is not identically zero and is invariant under $\hat{H}_{1}\subset\mathrm{GL}(V_{1})$. But Proposition 1.9 implies that there is no semi-invariant polynomial of a primitive group in dimension $3$ satisfying one of the listed conditions on degree and order.
Partition No. 14: $\pi=(4,2)$. Let $\dim(V_{1})=4$ and $\dim(V_{2})=2$. Consider 4.1 for $i=2$. If $\bar{H}_{2}$ is $A_{5}$, the kernel $\bar{K}$ is either all of $A_{5}$ or trivial. If it is trivial, $[H:Z(H)]$ is reduced by a factor of $60$ from its maximum value, and we’re done by Table 2. Otherwise, $d\geq 12$, which is greater than the max degree from the table.
Hence for an exceptional example we need $|\bar{H}_{2}|\leq 24$, and Theorem 3.4 gives $|G|\leq 24[\hat{H}_{1}:Z(\hat{H}_{1})]d^{2}$. Comparing this to the Fermat hypersurface, we need either $d=4,5$ and $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 7680$ or $d=3$ and $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 2430$. The polynomial $f(x_{0},x_{1},x_{2},x_{3},0,0)$, which is nonzero by Lemma 4.2, is an invariant of the primitive representation $\hat{H}_{1}\subset\mathrm{GL}(V_{1})$ and would have to satisfy the order and degree conditions above. Proposition 1.10 rules out the existence of such a polynomial.
Partition No. 22: $\pi=(4,3)$. Let $\dim(V_{1})=4$ and $\dim(V_{2})=3$. Consider 4.1 for $i=2$. Suppose first that $\bar{H}_{2}$ is $A_{6}$ or the Hessian group. The possible kernels $\bar{K}$ are $\bar{H}_{2}$ itself, the index $3$ normal subgroup of the Hessian, or a group of order at most $18$. (This is because the next largest normal subgroup of the Hessian group has order $18$.) In the latter scenario, the order of $H/Z(H)$ falls short of its maximum by a factor of at least $360/18=20$, which is greater than the value $7.62$ from Table 2. But if $\bar{K}$ is one of the other groups listed, $d\geq 6$, which is greater than the maximum of $4$ from the table. Hence the largest possible $\bar{H}_{2}$ is $\mathrm{PSL}_{2}(7)$.
Thus Theorem 3.4 gives $|G|\leq 168[\hat{H}_{1}:Z(\hat{H}_{1})]d^{2}$. Comparing this to the Fermat hypersurface, we get $d=3$ and $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 7290$. Using Lemma 4.2, the polynomial $f(x_{0},x_{1},x_{2},x_{3},0,0,0)$ is nonzero and is an invariant of the primitive representation $\hat{H}_{1}\subset\mathrm{GL}(V_{1})$. By Proposition 1.10, no group satisfies the necessary condition on order and degree.
Partition No. 23: $\pi=(4,2,1)$. Let $\dim(V_{1})=4,\dim(V_{2})=2$, and $\dim(V_{3})=1$. Consider the composition from 4.1 with $i=2$. If $\bar{H}_{2}$ is $A_{5}$, the kernel $\bar{K}$ is all of $A_{5}$ or trivial. If it is trivial, $[H:Z(H)]$ is reduced by a factor of $60$ from its maximum value, and we’re done by Table 2. Otherwise, $d\geq 12$, which is greater than the max degree from the table.
Hence for an exceptional example we need $|\bar{H}_{2}|\leq 24$, and Theorem 3.4 gives $|G|\leq 24[\hat{H}_{1}:Z(\hat{H}_{1})]d^{3}$. Comparing this to the Fermat hypersurface, this forces $d=3$ and $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 17010$. The polynomial $f(x_{0},x_{1},x_{2},x_{3},0,0,0)$, which is nonzero by Lemma 4.2, is an invariant of the primitive representation $\hat{H}_{1}\subset\mathrm{GL}(V_{1})$; this polynomial and $\hat{H}_{1}$ would have to satisfy the degree and order conditions above. Proposition 1.10 rules out the existence of such a polynomial.
Partition No. 30: $\pi=(4^{2})$. In this case, we have two primitive constituents $\hat{H}_{1}$ and $\hat{H}_{2}$, both primitive subgroups of $\mathrm{GL}_{4}({\mathbb{C}})$. Suppose one constituent, say $\hat{H}_{1}$ without loss of generality, satisfies $\bar{H}_{1}=\mathrm{PSp}_{4}(3)$ or $\bar{H}_{1}=N/Z(N)$, for the group $N$ from Proposition 1.10 (2). Then the kernel $\bar{K}$ of the composition in 4.1 for $i=1$ is a normal subgroup of $\bar{H}_{1}$, so that it is all of $\mathrm{PSp}_{4}(3)$ (in case (1)), either $N/Z(N)$ or its index $2$ subgroup (in case (2)), or has index at least $720$ in $\bar{H}_{1}$. When the index is at least $720$, $[H:Z(H)]$ falls short of the upper bound by at least that number, so we’re done by Table 2.
In the remaining cases, we use 4.1 to conclude $d\geq 8$, which is larger than the maximum degree value from the table. Hence $[\hat{H}_{i}:Z(\hat{H}_{i})]\leq 7200$ for each $i$. On the other hand, Theorem 3.4 gives
(3) $|G|\leq 2[\hat{H}_{1}:Z(\hat{H}_{1})][\hat{H}_{2}:Z(\hat{H}_{2})]d^{3}.$
Comparison to $B((1^{8}),d)$ yields $d=3$ and
$[\hat{H}_{1}:Z(\hat{H}_{1})][\hat{H}_{2}:Z(\hat{H}_{2})]\geq 14696640.$
The factor of $2$ in front of (3) comes from the possibility of a symmetry between the two factors, but this can only occur if $\hat{H}_{1}\cong\hat{H}_{2}$. Consulting the list in Proposition 1.10 gives that the only remaining possibility is that $\hat{H}_{i}/Z(\hat{H}_{i})=Q/Z(Q)$ for each $i$, where $Q$ has an index $2$ subgroup $2.A_{5}\times 2.A_{5}/Z$ and $Z$ is a central subgroup of order $2$. (Here $Q$ is the largest group from Proposition 1.10(3)).
The last thing to do is see that $d=3$ cannot occur for this choice of primitive constituents. Returning to 4.1 again for $i=1$, $\bar{K}$ is a normal subgroup of $Q/Z(Q)$ which is either the whole subgroup, a subgroup of index $2$, or trivial. If it is trivial, $[H:Z(H)]$ falls short of the upper bound by at least $7200$ and we’re done. In the other two cases, the conclusion of 4.1 gives that every term of $f$ has even degree in the variables $x_{0},x_{1},x_{2},x_{3}$ corresponding to $V_{1}$. But we can repeat this whole argument with 4.1 again for $i=2$ and get that every term of $f$ has even degree in $x_{4},x_{5},x_{6},x_{7}$. Hence $f$ has even degree, and $d\neq 3$, a contradiction.
Partition No. 31: $\pi=(4,2^{2})$. Let $\dim(V_{1})=4$ and $\dim(V_{2})=\dim(V_{3})=2$. We’ll work by finding stronger bounds on the orders of each primitive constituent. Suppose that at least one of the dimension $2$ constituents, say $\hat{H}_{2}$, is an extension of $A_{5}$. Then the group $\bar{K}$ in 4.1 for $i=2$ is trivial or all of $A_{5}$. If it is trivial, the upper bound on $[G:Z(G)]$ decreases by $60$ and we’re done by Table 2. If it is all of $A_{5}$, $d\geq 12$, contradicting the max degree from the table. Hence each of the two-dimensional constituents satisfies $[\hat{H}_{i}:Z(\hat{H}_{i})]\leq 24$. Now Theorem 3.4 gives
$|G|\leq 2\cdot 24^{2}[\hat{H}_{1}:Z(\hat{H}_{1})]d^{3}.$
Comparing to the Fermat hypersurface gives that $d=3$ and $[\hat{H}_{1}:Z(\hat{H}_{1})]\geq 8505$. Hence $\hat{H}_{1}$ must be as in (1) or (2) in Proposition 1.10 (in the latter case, $\hat{H}_{1}:Z(\hat{H}_{1})\cong N/Z(N)$ for the entire normalizer $N$, in particular). Now we consider 4.1 for $i=1$, so that $\bar{K}$ is a normal subgroup of $\bar{H}_{1}$. This must either be all of $\bar{H}_{1}$, have index $2$, or else its index is so large that $|H/Z(H)|$ is reduced by a factor of at least $720$, and we’d be done. In the former cases, we’d have $d\geq 8$, contradicting $d=3$. ∎
4.2. The Exceptional Examples
Putting together Corollaries 4.4, 4.5, 4.7, 4.9, 4.11, and Lemma 4.12, the only remaining exceptional partitions to consider are No. 1, No. 2, No. 4, No. 6, No. 13, and No. 18. We won’t consider No. 1, the partition $\pi=(2)$, because this would correspond to a zero-dimensional hypersurface $X$.
The remaining cases are partitions of $N=3,4$, or $6$, so in particular we’ve finished proving that no $X$ with corresponding partition $\pi\neq(1^{N})$ can match or exceed the Fermat unless $\dim(X)=n=1,2,4$. Each of the remaining partitions will actually yield hypersurfaces with larger automorphism group than the Fermat hypersurface in certain degrees. In this section, we’ll find which degrees can occur and prove that the examples from Section 2 are unique, up to isomorphism.
Proof of Theorem 2.3:.
We’ll consider each of the remaining partitions in turn. We begin each case by assuming $|G|\geq B((1^{N}),d)$.
Partition No. 2: $\pi=(3)$.
[…]
Hence any exceptional example would have $[G:Z(G)]d\geq 6d^{3}$, or $[G:Z(G)]\geq 6d^{2}$. In particular, $[G:Z(G)]\geq 6\cdot 3^{2}=54$. There are four possibilities, following Proposition 1.9.
[…]
If $G/Z(G)=H/Z(H)$ for a subgroup $H$ of the Hessian group with $[H:Z(H)]\geq 54$, $d\leq 6$ since $[H:Z(H)]\leq 216$. But the smallest degree semi-invariants also have degree $6$ in this case, so $d=6$. The largest possible $|\mathrm{Aut}(X)|$ is order $216$, so this falls short of the automorphism group of the Wiman sextic, and we get no new maximal example.
[…]
Partition No. 4: $\pi=(4)$.
[…]
Hence any exceptional example would have $[G:Z(G)]d\geq 24d^{4}$, or $[G:Z(G)]\geq 24d^{3}$. Since $d\geq 3$, in particular it’s true that $[G:Z(G)]\geq 648$, exactly the condition in Proposition 1.10. We need only check these cases.
We see that in cases (1), (4), and (5) of Proposition 1.10, that $[G:Z(G)]<24d^{3}$, where $d$ is the minimum degree of a semi-invariant. In case (3), when $1152<[G:Z(G)]\leq 7200$, $d\leq 6$ and there are possible semi-invariants in degrees $4$ or $6$, but they are all powers of the quadric $x_{0}x_{3}-x_{1}x_{2}$ and do not define a smooth surface. If $[G:Z(G)]\leq 1152$, $d<4$ so there are no available semi-invariants to define a surface.
Lastly, in case (2), $[G:Z(G)]\leq 11520$ so $d\leq 7$. We therefore can’t get any semi-invariants of the required degree unless $G/Z(G)\cong H/Z(H)$ for $H\subset N$ the subgroup corresponding to the exceptional embedding of $S_{5}\subset S_{6}$ (or its index $2$ subgroup). Either of these subgroups satisfies $[H:Z(H)]\leq 11520/6=1920$, so $d\leq 4$, and $d\leq 3$ for the index $2$ subgroup. By the conditions on semi-invariants from Proposition 1.10, we must have $d=4$, and the equation $f$ is the unique quartic from that proposition. Therefore, $X$ is isomorphic to Example 2.7.
Partition No. 6: $\pi=(2^{2})$.
In this case, the group $G=\mathrm{Lin}(f)$ is not primitive, but instead has two primitive constituents $\hat{H}_{1},\hat{H}_{2}$ of dimension $2$. The only possibilities for these constituents are central extensions of $A_{4}$, $S_{4}$, or $A_{5}$. By Theorem 3.4 we have
(4) $|G|\leq 2[\hat{H}_{1}:Z(\hat{H}_{1})][\hat{H}_{2}:Z(\hat{H}_{2})]d^{2}.$
On the other hand, we need $|G|\geq 24d^{4}$, so $[\hat{H}_{1}:Z(\hat{H}_{1})][\hat{H}_{2}:Z(\hat{H}_{2})]\geq 12d^{2}$. Suppose $\bar{H}_{1}=A_{5}$. Using 4.1 for $i=1$, the kernel $\bar{K}$ is trivial or all of $A_{5}$. If it is trivial, $|G|\leq 2[\hat{H}_{2}:Z(\hat{H}_{2})]d^{2}$ and we’d get $d<3$. Thus, the kernel is all of $A_{5}$ and $d\geq 12$. Only the choice $\bar{H}_{1}=\bar{H}_{2}=A_{5}$ gives a large enough group for this value of $d$, and each term of $f$ must symmetrically be an invariant polynomial in the variables $x_{2},x_{3}$ of the second summand as well. Finally, we know $d\leq 17$ from Table 2, so the only possibility is that $f$ is the sum of the unique degree $12$ invariant in both pairs of variables. Therefore, $X$ is isomorphic to Example 2.9.
If neither constituent is an extension of $A_{5}$, consider the case $\bar{H}_{1}=\bar{H}_{2}=S_{4}$. Using (4) again gives $d\leq 6$. We examine the possibilities for the subgroup $H/Z(H)\subset\bar{H}_{1}\times\bar{H}_{2}$. If it is the whole group, by 4.1 applied to either projection, $d=6$ and $f$ must be an invariant polynomial in each pair of variables, forcing it to be the sum of the unique degree $6$ invariant in both pairs of variables. Therefore, $X$ is isomorphic to Example 2.8.
If $H/Z(H)$ is a proper subgroup of the product, it is either not symmetric under the interchange of factors (so we remove the $2$ in the front of the bound in (4)) or has index at least $4$. The same reduction in order happens when we change the constituents: if $\bar{H}_{1}=A_{4}$ and $\bar{H}_{2}=S_{4}$, we must also remove the factor of $2$, and if $\bar{H}_{1}=\bar{H}_{2}=A_{4}$, then $[\hat{H}_{1}:Z(\hat{H}_{1})][\hat{H}_{2}:Z(\hat{H}_{2})]=144$. In all of the cases from this paragraph,
$|G|\leq 288Z(H)\leq 288d^{2}$
so $d=3$. But the oddness of the degree means $Z(H)$ is a proper subgroup of $Z(\hat{H})=Z(\hat{H}_{1})\times Z(\hat{H}_{2})$ (using Theorem 3.2), so in fact $|G|\leq 144d^{2}$. This is too small to compete with the Fermat for any degree $d\geq 3$.
Partition No. 13: $\pi=(6)$.
[…]
Partition No. 18: $\pi=(2^{3})$.
In this case, the group $G=\mathrm{Lin}(f)$ is not primitive, but instead has three primitive constituents $\hat{H}_{1},\hat{H}_{2},\hat{H}_{3}$ of dimension $2$. By Theorem 3.4 we have
(5) $|G|\leq 6[\hat{H}_{1}:Z(\hat{H}_{1})][\hat{H}_{2}:Z(\hat{H}_{2})][\hat{H}_{3}:Z(\hat{H}_{3})]d^{3}.$
In contrast, the Fermat equation has automorphism group of order $|G_{F}|=720d^{6}$. Even setting the max value of $60$ for each $[\hat{H}_{1}:Z(\hat{H}_{1})]$, we must have $d\leq 12$. First, suppose at least one primitive constituent $\hat{H}_{1}$ is a central extension of $A_{5}$. Then applying 4.1 for $i=1$, the kernel $\bar{K}$ is all of $A_{5}=\bar{H}_{1}$ or trivial. If it is trivial, this reduces the maximum possible order of $H/Z(H)$ by $60$ and the modified (5) gives $d\leq 3$. But if $d=3$, the degree being odd forces $|Z(H)|\leq\frac{d^{3}}{2}$, and $26\cdot 60^{2}\cdot d^{3}/$ 2 is smaller than $720d^{6}$, even for $d=3$.
So $\bar{K}$ is all of $A_{5}$, and $d=12$ is forced by 4.1. The only choice of primitive constituents consistent with this bound is $\bar{H}_{1}=\bar{H}_{2}=\bar{H}_{3}=A_{5}$. Since we can apply the same argument to each of the three factors, the only possible equation is the sum of the unique degree $12$ invariant polynomial in each pair of variables. Therefore, $X$ is isomorphic to Example 2.11.
If no primitive constituent is an extension of $A_{5}$, suppose $\bar{H}_{1}=S_{4}$. Unless the other two constituents are also extensions of $S_{4}$, the group $G/H$ cannot be a transitive subgroup of $S_{3}$ and so the coefficient $6$ in (5) is reduced to at most $2$. In this case, $d<3$, a contradiction. (We also get $d<3$ if all three constituents are extensions of $A_{4}$.) So $\bar{H}_{1}=\bar{H}_{2}=\bar{H}_{3}=S_{4}$ and the inequality in (5) gives $d\leq 4$.
Considering possible subgroups $H/Z(H)\subset\bar{H}_{1}\times\bar{H}_{2}\times\bar{H}_{3}$, either $H/Z(H)$ is this whole group, forcing $d\geq 6$ (a contradiction) or $H/Z(H)$ is a proper subgroup. This proper subgroup is either not fully symmetric under the permutation of factors, or it has index at least $8$. Either way, the bound on $|G|$ drops to the point that $d<3$, completing the proof.
Uniqueness for other $(n,d)$.
We’ve now dealt with all exceptional cases. In all other dimensions $n=N-2$ and degrees $d$ besides the seven pairs listed in Theorem 2.3, we’ve proven that the group $G=\mathrm{Lin}(f)$ has order $|G|<B((1^{N}),d)$ unless the partition $\pi$ of $G$ equals $(1^{N})$. Therefore, if a hypersurface $X_{d}=\{f=0\}\subset\mathbb{P}^{n+1}$ has automorphism group of order at least that of the Fermat hypersurface and $(n,d)$ is not one of the seven special pairs, then the associated $G=\mathrm{Lin}(f)$ has partition $\pi=(1^{N})$. Further, $|G|$ must achieve the upper bound $B((1^{N}),d)$. We’ll show under these assumptions that $X$ is isomorphic to the Fermat hypersurface.
Because $|G|$ already equals $B((1^{N}),d)$, applying Collins’ replacement theorem (Theorem 3.2) to produce a new group $\hat{G}$ outputs a group of the same order. Following the notation of that theorem, since $H\subset\hat{H}$, we must have $H=\hat{H}=\hat{H}_{1}\times\cdots\times\hat{H}_{N}$, and $G/H=S_{N}$ is the full symmetric group. Each $\hat{H}_{i}$ is a finite subgroup of $\mathrm{GL}(V_{i})\cong{\mathbb{C}}^{*}$ and acts trivially on all other $V_{i}$. This means that each term of $f$ must have degree a multiple of $|\hat{H}_{i}|$ in each variable $x_{i}$. The product of the $|\hat{H}_{i}|$ is $d^{N}$ and $f$ has degree $d$, so the only possibility is $\hat{H}_{i}=\mu_{d}$ for each $i$. Thus, $f$ is a linear combination of the monomials $x_{i}^{d}$. Finally, $f$ is invariant under the symmetric group permuting the subspaces $V_{i}$, so $f$ must in fact be the Fermat polynomial. ∎